### v1

# Geometric Origin of Quantum Entanglement

Abstract
We investigate massless representations related to the extension of Poincarè group constructed in [1]. These representations differ from Wigner’s ones of standard Poincarè group because the stabilizer of a lightlike momentum in the extended group is $ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}Z_{2}$, with factor $Z_{2}=\{1,-1\}$ generated by involution $\Lambda_{\infty}(\theta,\phi)$ which represents infinite velocity limit of a superluminal boost along spatial direction identified by polar and azimuthal angles $\theta,\phi$. The unitary irreducible representations (UIRs) of massless particles in this extension must decompose as a direct sum of a massless forward (positive zeroth component momentum) and massless backward (negative zeroth component momentum) Wigner’s representations linked by internal two valued degree of freedom given by the two possible eigenvalues of $U(\Lambda_{\infty})$. We prove that these representations are unitarily equivalent to entangled states of two qubits. This provides a geometric origin of quantum entanglement for photons in the framework of quantum field theory: photons appear as superpositions of backward and forward propagating electromagnetic waves depending on the eigenvalue of $U(\Lambda_{\infty})$ and this dependency gives rise to correlations between the values of local observables identical to those experienced with an entangled state of two qubits. Finally we describe an experiment capable of distinguishing the two eigenvalues of $U(\Lambda_{\infty})$ providing experimental falsification of the theory.
1 Introduction
Quantum entanglement is usually introduced as a feature of composite quantum systems: given a tensor product $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, there exist pure states that cannot be written as product vectors. In nonrelativistic quantum mechanics this structure is taken as primitive. By contrast, in relativistic quantum theory the state space is constrained by spacetime symmetry: one starts from unitary (or projective) representations of the Poincaré group and builds quantum fields and particles from them. It is then natural to ask whether some instances of entanglement may have a direct geometric origin in the representation theory of spacetime symmetries.
In [1] we constructed an extension of the proper orthochronous Lorentz group $\mathrm{SO}(3,1)^{+}_{\uparrow}$ that includes superluminal observers via involutive matrices $\Lambda_{\infty}(\theta,\phi)$ arising as infinite–velocity limits of superluminal boosts. The resulting extended group
$\mathcal{L}_{\mathrm{ext}}\;\cong\;\bigl(\mathrm{SO}(3,1)^{+}_{\uparrow}\rtimes_{\mathrm{Ad}_{\Lambda_{\infty}}}Z_{2}\bigr)\times Z_{2}$
shares the same identity component as the ordinary Lorentz group $O(3,1)$ but has disconnected components generated by $\Lambda_{\infty}(\theta,\phi)$ and $\Lambda_{-\infty}(\theta,\phi)$ in place of parity and time-reversal. When translations are included, this leads to an extended Poincaré group $\mathcal{P}_{\mathrm{ext}}=\mathbb{R}^{1,3}\rtimes\mathcal{L}_{\mathrm{ext}}$ whose unitary irreducible representations (UIRs) differ from the standard Wigner classification in two essential ways: (i) timelike and spacelike orbits are merged into a single tachyon/massive multiplet, and (ii) in the massless sector the stability subgroup is enlarged from Wigner’s $ISO(2)$ to
$ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}Z_{2},\qquad Z_{2}=\{1,-1\},\quad\Lambda_{-\infty}^{2}=\mathbb{I}.$
The presence of this extra $Z_{2}$ factor in the stabiliser of a lightlike orbit of the extended group has a direct and unavoidable representation–theoretic consequence. Every massless UIR of $\mathcal{P}_{\mathrm{ext}}$ is no longer a single forward Wigner representation, but comes as a two components object
$\pi_{\mathrm{ml}}^{\mathrm{ext}}\;\simeq\;\pi^{\text{fwd}}_{\varepsilon}\oplus\pi^{\text{bwd}}_{\varepsilon},\qquad\varepsilon=\pm 1,$
where $\pi^{\text{fwd}}$ and $\pi^{\text{vwd}}$ are forward and backward lightlike Wigner’s UIRs of ordinary Poincaré group, and $\varepsilon$ is the eigenvalue of the unitary operator $U(\Lambda_{\infty})$. From the extended group viewpoint, the massless sector always carries an additional binary internal label $\varepsilon=\pm 1$ attached to the pair $(\pi^{+},\pi^{-})$.
The central observation of the present paper is that this binary structure realises, in a spacetime–geometric way, the algebraic pattern usually associated with entangled qubit pairs. The representation space of $\pi_{\mathrm{ml}}^{\mathrm{ext}}$ can be written as a direct sum
$\mathcal{H}_{\oplus}=\mathcal{H}^{\text{fwd}}\oplus\mathcal{H}^{\text{bwd}},$
where the above summands carry forward and backward massless Wigner UIRs respectively, while the action of $U(\Lambda_{\infty})$ exchanges the two sectors up to a sign determined by $\varepsilon$. We show that there exist:
  • a unitary sector isometry
[…]
    where $\mathcal{H}$ is isomorphic to $\mathcal{H}^{\text{fwd}}\simeq\mathcal{H}^{\text{bwd}}$, and
  • a natural *–homomorphism of observable algebras
    $\iota:\ \mathcal{B}(\mathcal{H})\otimes M_{2}(\mathbb{C})\ \longrightarrow\ \mathcal{B}(\mathcal{H}_{\oplus}),$
such that for every state $\Psi\in\mathcal{H}_{\oplus}$ and for every local observable $A\otimes B$ one has
$\langle\Psi,\ \iota(A\otimes B)\,\Psi\rangle_{\mathcal{H}_{\oplus}}=\langle V\Psi,\ (A\otimes B)\,V\Psi\rangle_{\mathcal{H}\otimes\mathbb{C}^{2}}.$
In other words, a massless UIR of the extended Poincaré group is operationally indistinguishable, on all observables of the form $A\otimes B$, from an entangled state of two qubits where one of the qubits encodes the binary degree of freedom $\varepsilon$ associated with $U(\Lambda_{\infty})$. Entanglement in this setting is therefore not an additional structure put in by hand, but a consequence of the geometry of the extended Lorentz symmetry.
The aim of this paper is to make this equivalence precise and to explore its physical implications. We first recall the construction of the extended Lorentz and Poincaré groups and the structure of their massless UIRs. We then make explicit the isometry between the direct–sum representation $\mathcal{H}^{+}\oplus\mathcal{H}^{-}$ and the tensor–product representation $\mathcal{H}\otimes\mathbb{C}^{2}$, and we construct the associated *–homomorphism of local observable algebras. Finally, we outline a local–tomography experiment, within standard quantum optics and interferometry, that should detect the additional binary degree of freedom $\varepsilon$ and thus provide an empirical test of the extended Poincaré symmetry underlying this geometric origin of quantum entanglement.
2 UIRs of Extended Poincaré Group
In this section we briefy recall the mathematical construction leading to the extension of Poincarè group including superluminal observers.
The construction in [1] begins by considering superluminal boosts of velocity $v>c$ and taking the limit $|v|\to\infty$ while keeping the direction fixed. This yields an involutive Lorentz transformation $\Lambda_{\infty}(\theta,\phi)$ whose action in the $(t,\vec{x})$ coordinates is
$t\;\mapsto\;\hat{n}\cdot\vec{x},\qquad\hat{n}\cdot\vec{x}\;\mapsto\;t,\qquad\hat{n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta).$
Conjugation by $\Lambda_{\infty}$ is a nontrivial automorphism of the proper orthochronous group ${SO}(3,1)^{+}_{\uparrow}$ and generates the extension
$\mathcal{L}_{\mathrm{ext}}\cong\ ({SO}(3,1)^{+}_{\uparrow}\rtimes_{{\text{Ad}}_{\Lambda_{\infty}}}\mathbb{Z}_{2})\times\mathbb{Z}_{2}$
where $\mathbb{Z}_{2}$ is $\{$-1,1}\}, the nontrivial automorphism $\text{Ad}_{\Lambda_{\infty}}$ is conjugation by $\Lambda_{\infty}(\theta,\phi)$ and conjugation with matrices corresponding to different values $(\theta^{\prime},\phi^{\prime})$ produces equivalent group extensions. Including translations we defined the extended Poincarè group
$\mathcal{P}_{\text{ext}}\;=\;\mathcal{T}\rtimes\mathcal{L}_{\text{ext}}$ (1)
with multiplication
$(h,a)\cdot(h^{\prime},a^{\prime})\;=\;\bigl(hh^{\prime},\,a+ha^{\prime}\bigr),$
where $\mathcal{T}\cong\mathbb{R}^{4}$ is translation group, $h\in\mathcal{L}_{\text{ext}}$ acts on $\mathcal{T}$ by the standard linear action on $\mathbb{R}^{4}$.
Exploiting the fact that
$\mathcal{P}_{\text{ext}}=((\mathcal{T}\rtimes{SO}(3,1)^{+}_{\uparrow})\rtimes_{{\text{Ad}}_{\Lambda_{\infty}}}\mathbb{Z}_{2})\times\mathbb{Z}_{2}=(\mathcal{P}_{0}\rtimes_{{\text{Ad}}_{\Lambda_{\infty}}}\mathbb{Z}_{2})\times\mathbb{Z}_{2}$ (2)
we are able to classify unitary irreducible representations (UIRs) of the extended group $\mathcal{P}_{\text{ext}}$ from the induced action of the extension group:
$Z=\{I,-I,\Lambda_{\infty},-\Lambda_{\infty}\}$ (3)
on UIRs of Poincarè group $\mathcal{P}_{0}$. In order to explain this we call $N\equiv\mathcal{P}_{0}=(\mathcal{T}\rtimes SO(3,1)^{+}_{\uparrow})$ and define
N^={Wigner’s UIRs of $\hat{N}$$\{\text${Wigner's UIRs of N}\} (4)
An element $z\in Z$ acts on $N$ via automorphisms, namely its action on a given element of the Poincarè group $n=(h,a)\in N$ with $h\in SO(3,1)^{+}_{\uparrow}$:
$n^{\prime}=z\cdot n$ (5)
is such that it exists an automorphism of $N$, $\alpha_{z}(n)$ such that
$n^{\prime}=\alpha_{z}(n)$ (6)
where $\alpha_{z}$ is defined as:
$\alpha_{z}(h,a)=\big(zhz^{-1},\,z\!\cdot\!a\big),\qquad z\in Z$ (7)
This induces an action on the set of charachters (i.e. UIRs) of $\mathcal{P}_{0}$:
$z\cdot\pi(n)=\pi(\alpha_{z}(n))\;\;\;\forall n\in N$ (8)
The set $\hat{N}$ is constituted by UIRs of Poincarè group which have been classified by Wigner in [2]. UIRs of Poincarè group are classified by the value of the four momentum vector $p=p_{\mu}$ together with the value of the invariant lenght $|L_{s}\cdot p|^{2}$ associated to the orbit
$O_{p}=\{L_{s}\cdot p|L_{s}\in SO(3,1)^{+}_{\uparrow}\}$ (9)
$p$ is called the representative of the induced UIR of Poincarè group. Given choice of $p_{\mu}$ an $SO(3,1)^{+}_{\uparrow}$-orbit representative in Wigner’s classification, its invariant could be $p^{\mu}p_{\mu}\neq 0$ or $p^{\mu}p_{\mu}=0$. Choosing a non massless ($p^{\mu}p_{\mu}\neq 0$) representative $p_{\mu}$ we have:
  • Massive Representations: Unitary irreps of SO(3) for time-like orbits
    $\pi^{+}_{\text{mass}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}>0\;\;\;p_{0}>0\}$ (10)
    $\pi^{-}_{\text{mass}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}>0\;\;\;p_{0}<0\}$ (11)
  • Tachyonic Representations: Unitary irreps of SO(2,1) for space-like orbits
    $\pi_{\text{tach}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}<0\;\;\;\}$ (12)
Choosing a massless ($p^{\mu}p_{\mu}=0$) we have
  • Massless Representations: Unitary irreps of ISO(2) for light-like orbits
    $\pi_{\text{massless}}^{\text{fwd}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}=0\;\;p^{0}>0\}$ (13)
    $\pi_{\text{massless}}^{\text{bwd}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}=0\;\;p^{0}<0\}$ (14)
Now, with reference to the action in (8) define, given a representative $\pi\in\hat{N}$, its $Z$-orbit:
$O_{\pi}=\{z\pi\;\;|\;\;z\in Z\}$ (15)
and the stabilizers
$Z_{\pi}=\{z\in Z\;\;|\;\;z\pi(n)=\pi(n)\}$ (16)
where $n$ is in $N\equiv\mathcal{P}_{0}$.
We have two different possibilities [3].
If $Z_{\pi}=\{e\}\;\;\forall\pi(n)\;\in\;\hat{N}$ with $e$ identity element of $\pi$ then $\exists\;U(z)$ such that
$U(z)\pi(n)U(z)=\pi(z^{-1}nz)\;\;\;\forall z\in Z\;\;$ (17)
and the UIRs of $\mathcal{P}_{0}$ (namely the set $\hat{N}$) consist of UIRs of the extended group too. In this case UIRs which are inequivalent as representations of $\mathcal{P}_{0}$ become equivalent for $\mathcal{P}_{\text{ext}}$. This is due to the fact that orbits of $SO(3,1)^{+}_{\uparrow}$ in momentum space preserve $p^{\mu}p_{\mu}$ with its sign while in $\mathcal{L}_{\text{ext}}$ the sign can change.
If, on the contrary, $Z_{\pi}\neq\{e\}$ for some $\pi$ then its elements may give rise to inequivalent UIRs depending on the action of z $\in Z_{\pi}$ on $\pi\in\hat{N}$.
2.1 Non Massless Representations
For a given choice of momentum $p_{\mu}$, the possible non-lightlike $z$-orbits of the extended group are those with $p^{\mu}p_{\mu}\neq 0$ and give rise to the following UIRs of the Poincarè group $\mathcal{P}_{0}$:
$\hat{N}=\{\pi^{+}_{\text{mass}},\pi^{-}_{\text{mass}},\pi_{\text{tach}}\}$ (18)
The stabilizers of the corresponding orbits in (15) $Z_{\pi}={e}$, $\forall\pi\in\hat{N}$ and from (17)we have:
$U(\Lambda_{\infty})\pi^{+}_{\text{mass}}(n)U(\Lambda_{\infty})^{-1}=\pi^{+}_{\text{mass}}(\Lambda_{\infty}n\Lambda_{\infty}^{-1})=\pi_{\text{tach}}$ (19)
$U(\Lambda_{-\infty})\pi^{+}_{\text{mass}}(n)U(\Lambda_{-\infty})^{-1}=\pi^{-}_{\text{mass}}(\Lambda_{-\infty}n\Lambda_{-\infty}^{-1})=\pi_{\text{tach}}$ (20)
$U(-\mathbb{I})\pi^{+}_{\text{mass}}(n)U(-\mathbb{I})^{-1}=\pi^{+}_{\text{mass}}((-\mathbb{I})n(-\mathbb{I}))=\pi^{-}_{\text{mass}}$ (21)
We thus have that the set of UIRs in (18) is a set of equivalent irreducible representations of the extended group and in order to specify the full UIR we need to derive the action of the operators $\bar{U}(z)$ for $z$ in $Z$ on wavefunctions. Again the trick is to start from the known Wigner’s representations for $z=I$ and derive the actions of $\bar{U}(z)$ for $z\neq I$ from the action of the extension group $Z$ on Wigner’s representations space. In [1] it is showed:
$\bar{U}(\Lambda_{\infty})\psi](p)=\psi(\Lambda_{\infty}^{-1}p)$ (22)
and similarly for $\bar{U}(-I)$ and $\bar{U}(-\Lambda_{\infty})$we have:
$\bar{U}(-I)\psi](p)=\psi(-p)\;\;\;\;[\bar{U}(-\Lambda_{\infty})\psi](p)=\psi((-\Lambda_{\infty})^{-1}p)$ (23)
2.2 Massless Representations
In [1] it is considered the lightlike zz–orbit
$\mathcal{O}_{0}\;=\;\{\,Lp_{0}\;|\;L\in SO(3,1)^{+}_{\uparrow},\ p_{0}^{2}=0\,\},$ (24)
with the standard representative chosen as
$p_{0}=(\omega,\omega\sin\theta\cos\phi,\omega\sin\theta\sin\phi,\omega\cos\theta),\qquad\omega>0.$ (25)
The usual Wigner classification for the ordinary Poincaré group $\mathcal{P}_{0}$ yields two massless UIRs, denoted
$\pi^{\text{fwd}},\ \pi^{\text{bwd}},$
corresponding to forward and backward light–cones ($p_{0}^{0}>0$ and $p_{0}^{0}<0$). They are both induced from unitary irreps of the Euclidean group $ISO(2)$ (helicity or continuous–spin class) and are inequivalent as representations of $\mathcal{P}_{0}$.
In the extended Lorentz group $\mathcal{L}_{\mathrm{ext}}$ the matrix $\Lambda_{-\infty}=-\Lambda_{\infty}$ satisfies $\Lambda_{-\infty}p_{0}=p_{0},$ hence it lies in the stability subgroup of the chosen representative. Consequently, the geometric little group at $p_{0}$ in $\mathcal{L}_{\mathrm{ext}}$ is
$ISO(2)\rtimes\{I,\Lambda_{-\infty}\}.$ (26)
Let $U_{0}$ be a unitary representation of the little group $ISO(2)$ associated with either helicity or continuous–spin. The induced Poincaré action on wavefunctions $\psi:\mathcal{O}_{0}\to\mathcal{H}_{0}$ has the usual form
$\bar{U}(a,h)\psi\,](p)\;=\;e^{\,i\,p\cdot a}\;U_{0}\!\big(s_{0}(h,p)\big)\;\psi(h^{-1}p),\qquad(a,h)\in\mathcal{P}_{0},$ (27)
and this construction gives the standard massless Wigner UIRs $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ depending on the choice of representative on the forward or backward orbit.
In contrast to $\Lambda_{-\infty}$, the transformation $\Lambda_{\infty}$ maps $p_{0}$ to the backward representative
$\Lambda_{\infty}(\theta,\phi)\,p_{0}=-p_{0},$ (28)
thus interchanging the forward and backward light–cones. In particular $\Lambda_{\infty}$ does not belong to the geometric stabilizer of $p_{0}$. However, at the level of massless UIRs of $\mathcal{P}_{\text{ext}}$, $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ are equivalent since the unitary operator representing $U(\Lambda_{\infty})$ is thus an intertwiner between $\pi_{\text{fwd}}$ and $\pi_{\text{bwd}}$
$U(\Lambda_{\infty})\;\pi^{\text{fwd}}(n)\;U(\Lambda_{\infty})^{-1}\;=\;\pi^{\text{fwd}}\bigl(\Lambda_{\infty}^{-1}n\Lambda_{\infty}\bigr)\;=\;\pi^{\text{bwd}}(n),\qquad\forall n\in\mathcal{P}_{0}.$ (29)
Thus, under the action of the discrete factor $Z_{(\theta,\phi)}=\{I,-I,\Lambda_{\infty},-\Lambda_{\infty}\}$ on the dual space $\widehat{\mathcal{P}_{0}}$, (i.e. the set of Wigner’s UIRs) the set $\Pi\equiv\{\pi^{\text{fwd}},\pi^{\text{bwd}}\}$ constitute a single equivalence class of UIRs in which the transformation $U(\Lambda_{\infty})$ is a non trivial stabilizer, namely:
$\Lambda_{\infty}\Pi=\Pi$ (30)
As a consequence, the induced massless UIRs of the extended Poincaré group $\mathcal{P}_{\mathrm{ext}}$ do not live on a single copy of $\mathcal{H}_{0}$, but rather on the direct sum
$\mathcal{H}_{\oplus}\;=\;\mathcal{H}_{\text{fwd}}\oplus\mathcal{H}_{\text{bwd}},$ (31)
where $\mathcal{H}_{\text{fwd}}$ and $\mathcal{H}_{\text{bwd}}$ carry $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$, respectively. Writing $\Psi=(\psi_{\text{fwd}},\psi_{\text{bwd}})\in\mathcal{H}_{\oplus}$, the extended action of $(a,h)\in\mathcal{P}_{0}$ is block–diagonal:
$\bar{U}(a,h)\Psi\,](p)=\bigl([\bar{U}_{\text{fwd}}(a,h)\psi_{\text{fwd}}](p),[\bar{U}_{\text{bwd}}(a,h)\psi_{\text{bwd}}](p)\bigr),$ (32)
with each component governed by (27).
The discrete elements of $\mathcal{L}_{\mathrm{ext}}$ act as follows. For $\Lambda_{\infty}$ we choose an operator $\bar{U}(\Lambda_{\infty})$ on $\mathcal{H}_{\oplus}$ of the form
$\bar{U}(\Lambda_{\infty})\Psi\,](p):=\bigl(C\,\psi_{\text{bwd}}(\Lambda_{\infty}^{-1}p),\;C^{-1}\,\psi_{\text{fwd}}(\Lambda_{\infty}^{-1}p)\bigr),$ (33)
where $C$ implements the intertwining between $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ as above. One checks, using a section $k_{0}$ adapted as
$k_{0}(\Lambda_{\infty}^{-1}p)\;=\;\Lambda_{\infty}^{-1}\,k_{0}(p),$ (34)
and the covariance relation
$s_{0}\!\big(\Lambda_{\infty}h\Lambda_{\infty}^{-1},\,p\big)\;=\;\Lambda_{\infty}\,s_{0}\!\big(h,\Lambda_{\infty}^{-1}p\big)\,\Lambda_{\infty}^{-1},$ (35)
that
$\bar{U}(\Lambda_{\infty})\,\bar{U}(a,h)\,\bar{U}(\Lambda_{\infty})^{-1}\;=\;\bar{U}\!\big(\Lambda_{\infty}a,\ \Lambda_{\infty}h\Lambda_{\infty}^{-1}\big),$ (36)
so $\bar{U}(\Lambda_{\infty})$ correctly represents the extension.
Similarly, for I-I we may choose
$\bar{U}(-I)\Psi\,](p):=\bigl(\psi_{\text{bwd}}(-p),\,\psi_{\text{fwd}}(-p)\bigr),$ (37)
which exchanges forward and backward sectors and satisfies
$\bar{U}(-I)\,\bar{U}(a,h)\,\bar{U}(-I)^{-1}\;=\;\bar{U}\!\big(-a,\ (-I)h(-I)^{-1}\big).$ (38)
The only further constraint on $\bar{U}(\Lambda_{\infty})$ is that $\bar{U}(\Lambda_{\infty})^{2}=\mathbb{I}.$ This restricts $C$ in (33) to a unitary with $C^{2}=\mathbb{I}$ on $\mathcal{H}_{0}$, so that its eigenvalues are $\pm 1$. Different signs for the eigenvalues of $C$ lead to two inequivalent massless UIRs of the extended group for each standard Wigner class (helicity or continuous–spin). In this sense, the lightlike orbit supports a doublet representation
$\pi^{\mathrm{ext}}_{\varepsilon}\;\sim\;\pi^{\text{fwd}}_{\varepsilon}\ \oplus\ \pi^{\text{bwd}}_{\varepsilon},\qquad\varepsilon=\pm 1,$
where $\varepsilon$ labels the choice of sign in the representation of the extended little group of lightlike momenta representatives.
3 Equivalence of Massless UIRs with Entangled Two–Qubits States
The two states:
$\Psi_{\varepsilon=\pm 1}=\frac{1}{\sqrt{2}}(\psi_{\text{fwd}}\oplus\pm\psi_{\text{bwd}})$ (39)
are eigenstates of $\bar{U}(\Lambda_{\infty})$ corresponding to eigenvalues $\pm 1$. We now choose an isomorphism $\mathcal{H}\simeq\mathcal{H}_{\text{fwd}}\simeq\mathcal{H}_{\text{bwd}}$, a set of orthonormal vectors $|\text{fwd}\rangle,|\text{bwd}\rangle$ spanning a two–dimensional complex space $\mathcal{K}\simeq\mathbb{C}^{2}$ and define the sector isometry in the represenation space (31):
$V:\mathcal{H}_{\oplus}\to\mathcal{H}\otimes\mathcal{K},\qquad V(\Psi_{\varepsilon=\pm 1})=\psi_{\text{fwd}}\otimes|\text{fwd}\rangle\pm\psi_{\text{bwd}}\otimes|\text{bwd}\rangle.$ (40)
This is unitary and preserves inner products.
Let $\mathcal{B}(\mathcal{H})$ be the algebra of bounded operators on $\mathcal{H}$ and $\mathcal{B}(\mathcal{K})$ the algebra of all $2\times 2$ matrices acting on the internal qubit $\text{span}\{|\text{fwd}\rangle,|\text{bwd}\rangle\}$. For any $A\in\mathcal{B}(\mathcal{H})$ and $B=[b_{ij}]\in\mathcal{B}(\mathcal{K})$ define the unital *–homomorphism
$\iota(A\otimes B)=\begin{pmatrix}b_{00}A&b_{01}A\\ b_{10}A&b_{11}A\end{pmatrix},\qquad\iota:\mathcal{B}(\mathcal{H})\otimes B(\mathcal{K})\;\to\;\mathcal{B}(\mathcal{H}_{\oplus}).$ (41)
It preserves products, adjoints, and the identity operator.
A direct calculation shows that for any linear superposition $\Psi$ of states in (39) we have
$\big\langle\Psi,\,\iota(A\otimes B)\,\Psi\big\rangle=\big\langle V\Psi,\,(A\otimes B)\,V\Psi\big\rangle.$ (42)
Therefore all observable predictions of the $\mathcal{P}_{\mathrm{ext}}$ massless UIR coincide with those of an entangled state of two qubits. The states in (39) correspond under $V$ to Bell-like states.
Thus entanglement is not added by hand but emerges from the geometry of the extended Lorentz symmetry.
In block matrix form (with respect to the decomposition $\mathcal{H}\oplus\mathcal{H}$) the action of $U(\Lambda_{\infty})$ in (33) is
$U(\Lambda_{\infty})=\begin{pmatrix}0&I_{\mathcal{H}}\\ I_{\mathcal{H}}&0\end{pmatrix}$, (43)
On the tensor product space we consider the action of $I\otimes\sigma_{x}$ where $\sigma_{x}$ is the Pauli matrix acting on $\mathcal{K}\equiv\text{span}\{|\text{fwd}\rangle,|\text{bwd}\rangle\}$ as:
$\sigma_{x}|\text{fwd}\rangle=|\text{bwd}\rangle,\quad\sigma_{x}|\text{bwd}\rangle=|\text{fwd}\rangle.$
and thus, in this basis, has the form:
$\sigma_{x}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, (44)
We now compare the action of $I\otimes\sigma_{x}$ on $V(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})$ with the action of $V$ on $U(\Lambda_{\infty})(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})$.
The first is:
$(I\otimes\sigma_{x})\,V(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})=(I\otimes\sigma_{x})\big(\psi_{\text{fwd}}\otimes|\text{bwd}\rangle+\psi_{\text{bwd}}\otimes|\text{fwd}\rangle\big)$ (45)
On the other hand,
$U(\Lambda_{\infty})(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})=\psi_{\text{bwd}}\oplus\psi_{\text{fwd}},$ (46)
so
$V\big(U(\Lambda_{\infty})(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})\big)=V(\psi_{\text{bwd}}\oplus\psi_{\text{fwd}})=\psi_{\text{bwd}}\otimes|\text{fwd}\rangle+\psi_{\text{fwd}}\otimes|\text{bwd}\rangle.$ (47)
Therefore
$(I\otimes\sigma_{x})\,V(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})\;=\;V\big(U(\Lambda_{\infty})(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})\big)$ (48)
Since this must hold for all wavefunctions in $\mathcal{H}_{\oplus}$ we conclude that, via the isometric identification $V$, the sector–swap operator $U(\Lambda_{\infty})$ on $\mathcal{H}\oplus\mathcal{H}$ is unitarily equivalent to the operator $\mathbf{1}\otimes\sigma_{x}$ on $\mathcal{H}\otimes\mathbb{C}^{2}$.
4 Experimental Proposal: Tomography of $\varepsilon$ Eigenvalue
In this section we clarify the operational meaning of the additional binary degree of freedom $\varepsilon=\pm 1$ associated with $U(\Lambda_{\infty})$, and we outline a concrete single–photon interferometric implementation in standard quantum optics. The goal is to connect the abstract decomposition
$\mathcal{H}_{\mathrm{ml}}^{\mathrm{ext}}\simeq\mathcal{H}_{\text{fwd}}\oplus\mathcal{H}_{\text{bwd}}$
with experimentally accessible photon modes.
4.1 Specialisation to a fixed direction and single–mode photons
In the extended Poincaré picture the extra structure in the massless sector comes from the choice of a spatial direction
$\hat{n}(\theta,\phi)=(\sin\theta\cos\phi,\ \sin\theta\sin\phi,\ \cos\theta),$
[…]
  • Choice of $\hat{n}$. Fixing $(\theta,\phi)$ amounts to choosing a physical propagation axis in space. In the proposed single–photon experiment we specialise, without loss of generality, to $\hat{n}=(0,0,1)$, i.e. the optical axis is aligned along the zz–direction. The interferometer arms then realise the two counterpropagating lightlike modes with momenta
    $p_{+}=\omega(1,0,0,1),\qquad p_{-}=\omega(1,0,0,-1),$
    which are the forward and backward representatives associated with this choice of $\hat{n}$. Rotating the entire setup in space would simply correspond to choosing a different pair $(\theta,\phi)$ in the abstract construction.
so $p_{0}=\omega(1,0,0,1),$ with $\Lambda_{\infty}=\Lambda_{\infty}(0,0)$ an involution exchanging the time coordinate $t$ with the spatial coordinate $z$. we can represent a fixed helicity component by kets
$|p_{+}\rangle\in\mathcal{H}_{\text{fwd}},\qquad|p_{-}\rangle\in\mathcal{H}_{\text{bwd}},$ (49)
corresponding to the two lightlike directions along +z+z and z-z. Under the isomorphism $\mathcal{H}_{\text{fwd}}\simeq\mathcal{H}_{\text{bwd}}\simeq\mathcal{H}$, these are mapped to the “direction basis”
$|+\rangle,\ |$-$\rangle\in\mathcal{H}$
Once the spatial direction $\hat{n}$ is fixed, the remaining geometric degree of freedom in the theory is the eigenvalue $\varepsilon=\pm 1$ of $U(\Lambda_{\infty})$ acting on the two–component massless UIR. In the optical realisation this binary label is encoded in the relative phase between the forward and backward components of the photon state.
Restricting to a single momentum mode and a fixed helicity, and using the identifications
[…]
$|\text{fwd}\rangle\equiv|H\rangle,\qquad|\text{bwd}\rangle\equiv|V\rangle,$
the prepared single–photon state at the output of the interferometer (with a controllable phase shift in one arm) can be written as
$|\Psi(\varphi)\rangle=\frac{1}{\sqrt{2}}\Big(|+\rangle\otimes|H\rangle+e^{i\varphi}\,|-\rangle\otimes|V\rangle\Big),$ (50)
where $\varphi$ is the experimentally tunable interferometric phase (e.g. adjusted with a phase shifter or by varying the optical path length in one arm).
In the abstract representation–theoretic description, the $U(\Lambda_{\infty})$–eigenstates are precisely the symmetric and antisymmetric superpositions
$|\Psi_{\varepsilon=\pm 1}\rangle=\frac{1}{\sqrt{2}}\Big(|+\rangle\otimes|H\rangle\pm|-\rangle\otimes|V\rangle\Big),$ (51)
which are obtained from (50) by setting
$\varphi=\begin{cases}0\mod 2\pi,&\varepsilon=+1,\\[3.0pt] \pi\mod 2\pi,&\varepsilon=-1.\end{cases}$ (52)
Thus the theory predicts that controlling the eigenvalue sector $\varepsilon$ is operationally equivalent to controlling the interferometric phase $\varphi$ between the two counterpropagating components.
4.2 Preparation stage
A preparation scheme consistent with the above identification is:
  1. 1.
    Use a heralded single–photon source to prepare photons in a well-defined wave packet peaked around the four–momentum $p_{0}=\omega(1,0,0,1)$, with fixed helicity (e.g. right–handed circular polarisation).
  2. 2.
    Send the photon into an interferometer (e.g. a Michelson or Mach–Zehnder configuration) aligned along the zz–axis, with a highly reflecting mirror in one arm so that the output state contains a coherent superposition of forward and backward propagation along $\pm\hat{z}$. At the level of the abstract representation, this realises a superposition of the two sector basis vectors $|+\rangle$ and $|$-$\rangle$.
  3. 3.
    Insert polarisation optics in the two arms so that the component travelling in the $\hat{z}$ direction emerges with polarisation $|H\rangle$ and the component travelling in the $\hat{z}$ direction emerges with polarisation $|V\rangle$. Relative phases between the two paths can be tuned with phase shifters to realise
    $\frac{1}{\sqrt{2}}\bigl(|+\rangle\otimes|H\rangle+\varepsilon\,|-\rangle\otimes|V\rangle\bigr),$
    with $\varepsilon=\pm 1$ determined by the interferometric phase (see 52).
From the viewpoint of the extended Poincaré group, the two classes of preparations with relative phase $\varepsilon=\pm 1$ are interpreted as populating the two eigenvalue sectors of $U(\Lambda_{\infty})$ associated with the lightlike orbit in the direction $\hat{n}=(0,0,1)$.
4.3 Measurement stage and dependence on $\varepsilon$
To detect $\varepsilon$ we must measure an observable whose effective action, in the abstract description, is $\iota(A\otimes\sigma_{x})$ (see (41)) with an $A$ that mixes the forward/backward sectors. A natural choice is the Pauli operator
$A=\sigma_{x}^{(\mathrm{dir})}=|+\rangle\langle-|+|-\rangle\langle+|$
on the direction subspace. On the polarization subspace (encoding the basis $|\text{fwd}\rangle,|\text{bwd}\rangle$) we are forced to measure $\sigma_{x}^{(\mathrm{pol})}$. In the two–qubit picture the corresponding correlation observable is
$O_{XX}=\sigma_{x}^{(\mathrm{dir})}\otimes\sigma_{x}^{(\mathrm{pol})}.$
A straightforward calculation in the basis
$\{|+\rangle\otimes|H\rangle,\ |+\rangle\otimes|V\rangle,\ |-\rangle\otimes|H\rangle,\ |-\rangle\otimes|V\rangle\}$
yields, for the state (51),
$\langle\Psi_{\varepsilon}|\,O_{XX}\,|\Psi_{\varepsilon}\rangle=\varepsilon.$ (53)
Thus the sign of the correlation between (i) direction measured in the XX–basis of the forward/backward sectors and (ii) polarisation measured in the XX–basis of the internal qubit is exactly the eigenvalue $\varepsilon$ of $U(\Lambda_{\infty})$.
In the optical realisation, $O_{XX}$ can be implemented with standard components:
  • A 50:50 beam splitter acting on the counterpropagating modes along $\pm\hat{z}$ interferes $|+\rangle$ and $|$-$\rangle$ into two output ports corresponding to the eigenstates of $\sigma_{x}^{(\mathrm{dir})}$.
  • In each output port a half–wave plate at $45^{\circ}$ followed by a polarising beam splitter measures polarisation in the $\{|+\rangle_{X},|-\rangle_{X}\}$ basis (eigenbasis of $\sigma_{x}^{(\mathrm{pol})}$).
  • Single–photon detectors at the exits record joint outcomes $(x_{\mathrm{dir}},x_{\mathrm{pol}})\in\{\pm 1\}\times\{\pm 1\}$, from which one reconstructs the correlation
    $E_{XX}=\sum_{x_{\mathrm{dir}},x_{\mathrm{pol}}=\pm 1}x_{\mathrm{dir}}x_{\mathrm{pol}}\,P(x_{\mathrm{dir}},x_{\mathrm{pol}}),$
    which in the idealised extended theory reproduces (53).
4.4 Interpretation and falsifiability
From the viewpoint of standard quantum optics, the above is a textbook single–photon entanglement experiment between a path–like degree of freedom (here encoding the two sectors associated with the lightlike direction $\hat{n}$) and polarisation. What is nontrivial in the present work is the representation–theoretic interpretation: the same setup constitutes a local tomography of the binary degree of freedom $\varepsilon$ associated with $U(\Lambda_{\infty})$ in the massless UIR of $\mathcal{P}_{\mathrm{ext}}$.
The extended theory predicts that:
  • There exist preparations and measurements, compatible with the extended symmetry, in which the sign of the correlation $E_{XX}$ can be associated with the eigenvalue $\varepsilon$ of $U(\Lambda_{\infty})$ for a fixed lightlike direction $\hat{n}$, and in which switching sector corresponds to a flip $E_{XX}\to-E_{XX}$ without changing local marginals on $\mathcal{H}_{+}$ or $\mathcal{H}_{-}$.
  • If, in all such optical realisations, only one effective value of $\varepsilon$ is observed (e.g. $E_{XX}$ never changes sign in regimes where the extended theory predicts that both sectors should be accessible), then the massless sector of the extended Poincaré group would be empirically constrained or falsified.
In practice, imperfections such as losses, detector inefficiency and mode mismatch will reduce $|E_{XX}|$ below unity; one would then look for a statistically significant change of sign of the reconstructed correlation as a function of controlled preparation parameters. Nevertheless, the conceptual link is clear: by fixing the direction $\hat{n}$ of $\Lambda_{\infty}$, identifying the forward/backward sectors with counterpropagating photon modes along $\pm\hat{n}$, and performing local tomography on the resulting two–qubit system, one can in principle probe the geometric degree of freedom $\varepsilon$ and thus test the proposed geometric origin of entanglement for massless fields.
4.5 Relation to existing single–particle entanglement experiments
The state and measurement we propose in this section are, from the point of view of standard quantum optics, instances of single–particle entanglement between two internal degrees of freedom. In particular, the Bell–like state
$|\Psi_{\varepsilon}\rangle\;\simeq\;\frac{1}{\sqrt{2}}\bigl(|+\rangle\otimes|H\rangle+\varepsilon\,|-\rangle\otimes|V\rangle\bigr),\qquad\varepsilon=\pm 1,$ (54)
is a special case of the generic path–polarisation entangled state of a single photon, where $|+\rangle,|$-$\rangle$ encode two spatial modes and $|H\rangle,|V\rangle$ encode two polarisation modes. Correlations of Pauli–type observables such as $\sigma_{x}^{(\mathrm{dir})}\otimes\sigma_{x}^{(\mathrm{pol})}$ have been measured in several experiments.
For example, Fiorentino et al. implement a deterministic controlled–NOT gate acting on two qubits carried by a single photon: one qubit is encoded in polarisation, the other in the spatial mode (momentum) of the photon [4]. Their preparation and analysis stages generate and characterise states of the form
$|\Phi(\phi)\rangle=\frac{1}{\sqrt{2}}\bigl(|0\rangle_{\mathrm{path}}\otimes|H\rangle+e^{i\phi}\,|1\rangle_{\mathrm{path}}\otimes|V\rangle\bigr),$ (55)
which are mathematically equivalent to our $|\Psi_{\varepsilon}\rangle$ for phases $\phi=0,\pi$ and with an appropriate identification of $|0\rangle_{\mathrm{path}},|1\rangle_{\mathrm{path}}$ with $|+\rangle,|$-$\rangle$. Full two–qubit tomography in mutually unbiased bases (including the $X$ basis on each qubit) is performed, so that correlations proportional to $\langle\sigma_{x}^{(\mathrm{path})}\otimes\sigma_{x}^{(\mathrm{pol})}\rangle$ are effectively accessed [4].
Similarly, Bera et al. propose a protocol in which intra–photon entanglement between path and polarisation is swapped to inter–photon entanglement using linear optics [5]. Their initial single–photon resource is again a maximally entangled state between path and polarisation, structurally identical to $|\Psi_{\varepsilon}\rangle$, and the protocol requires the ability to prepare, transform and analyse such states in various local bases.
[…]
From the purely operational point of view, these experiments show that
  • a state preparation equivalent (up to local unitaries) to $|\Psi_{\varepsilon}\rangle$ is experimentally standard,
  • joint measurements of Pauli observables on the two internal degrees of freedom—including $X\otimes X$–type correlations—are feasible and have been performed, and
  • the observed correlations agree with the quantum predictions for a maximally entangled two–qubit state within experimental accuracy.
The novelty of the present work does not lie in the optical technology required to prepare and measure such single–photon two–qubit states, but in the representation–theoretic interpretation. In the above experiments, the two qubits are treated as abstract, independent Hilbert–space degrees of freedom (path, polarisation, spin), and their entanglement is a kinematical feature of the chosen encoding. In our framework, by contrast,
  1. 1.
    the “path” degree of freedom $|+\rangle,|$-$\rangle$ is identified with the forward/backward lightlike sectors associated with a fixed null momentum orbit of the extended Poincaré group, and
  2. 2.
    the second qubit corresponds to the binary internal label $\varepsilon=\pm 1$ associated with the eigenvalues of the superluminal involution $U(\Lambda_{\infty})$ in the massless UIR of $\mathcal{P}_{\mathrm{ext}}$.
In other words, existing experiments already implement the same mathematical structure—a two–qubit Bell state and local Pauli measurements—but they do not test the geometric claim that this structure arises from the extended Lorentz symmetry and that the internal qubit is nothing but the eigenvalue sector of $U(\Lambda_{\infty})$. Our proposed experiment is therefore best viewed as a reinterpretation and a targeted adaptation of standard single–photon entanglement setups, designed to perform a tomography of the $\varepsilon$ degree of freedom tied to the extended Poincaré representation, rather than a test of quantum mechanics itself.
5 Conclusion
We have shown that the massless representations of the extended Poincaré group necessarily possess a two–component internal structure correlating forward/backward Wigner sectors. This structure is mathematically equivalent, under local observables, to the entanglement of two qubits. Therefore quantum entanglement has a geometric origin in the additional discrete superluminal symmetry generated by $\Lambda_{\infty}$.
A quantum state tomography experiment doable within standard quantum optics can in principle distinguish the two eigenvalues of $U(\Lambda_{\infty})$ giving experimental evidence of the representation-theoretic structure of massless particles in the extended Poincarè group which are intrinsically entangled objects.
[…]
Figure 1: Schematic of the proposed single–photon experiment. BS1 prepares a superposition of counterpropagating modes $|+\rangle,|$-$\rangle$ (forward/backward sectors). The elements $H,V$ encode polarisation $|H\rangle,|V\rangle$ on each arm, implementing the state $\tfrac{1}{\sqrt{2}}(|+\rangle\otimes|H\rangle+\varepsilon|-\rangle\otimes|V\rangle)$. BS2 measures $\sigma_{x}$ on the direction qubit; in each output, a HWP (Half Wave Plate) + PBS (Polarization Beam Splitter) measures $\sigma_{x}$ on polarisation. Joint detector clicks ($D_{i}$) estimate $\langle\sigma_{x}^{(\text{dir})}\otimes\sigma_{x}^{(\text{pol})}\rangle=\varepsilon$.