### v2

# Geometric Quantum Computation

Abstract
Massless unitary irreducible representations of the extension of Poincarè group constructed in [1] differ from Wigner’s ones of standard Poincarè group because the stabilizer of a lightlike momentum in the extended group is $ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}\mathbb{Z}_{2}$, with factor $\mathbb{Z}_{2}=\{1,-1\}$ generated by involution $\Lambda_{\infty}(\theta,\phi)$ which represents infinite velocity limit of a superluminal boost along spatial direction identified by polar and azimuthal angles $\theta,\phi$. These must decompose as a direct sum of a massless forward (positive zeroth component momentum) and massless backward (negative zeroth component momentum) Wigner’s representations linked by internal two valued degree of freedom given by the two possible eigenvalues of $U(\Lambda_{\infty})$. We prove that these representations are operationally equivalent (on local observables) to entangled states of a two-qubit system. This provides a geometric origin of quantum entanglement for photons in the framework of quantum field theory: photons appear as superpositions of backward and forward propagating electromagnetic waves depending on the eigenvalue of $U(\Lambda_{\infty})$ and this dependency gives rise to correlations between the values of local observables identical to those experienced with an entangled state of two qubits. We then describe an experiment capable of distinguishing the two eigenvalues of $U(\Lambda_{\infty})$ providing experimental falsification of the theory. Using the above experiment as a primitive we introduce a model of quantum computation in which the logical qubit is not postulated as a two level system, but arises as the representation space of quantum states describing the entanglement between degrees of freedom of a single photon (e.g. helicity and direction of propagation). We call the resource of the computation single photon entanglement qubit and show it is canonically isomorphic to a standard qubit: its pure states form a Bloch sphere, its logical observables close an $\mathfrak{su}(2)$ algebra, and its admissible physical transformations act as automorphisms of the Bloch sphere. We then construct single photon entanglement qubits entangling gate exploiting parity measurements of physical degrees of freedom. Finally, we establish universality of the computational model in the conventional circuit sense.
1 Introduction
Quantum entanglement is usually introduced as a feature of composite quantum systems: given a tensor product $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, there exist pure states that cannot be written as product vectors. In nonrelativistic quantum mechanics this structure is taken as primitive. By contrast, in relativistic quantum theory the state space is constrained by spacetime symmetry: one starts from unitary representations of the Poincaré group and builds quantum fields and particles from them. It is then natural to ask whether some instances of entanglement may have a direct geometric origin in the representation theory of spacetime symmetries.
In [1] we constructed an extension of the proper orthochronous Lorentz group $\mathrm{SO}(3,1)^{+}_{\uparrow}$ that includes superluminal observers via involutive matrices $\Lambda_{\infty}(\theta,\phi)$ arising as infinite–velocity limits of superluminal boosts. The resulting extended group
$\mathcal{L}_{\mathrm{ext}}\;\cong\;\bigl(\mathrm{SO}(3,1)^{+}_{\uparrow}\rtimes_{\mathrm{Ad}_{\Lambda_{\infty}}}\mathbb{Z}_{2}\bigr)\times\mathbb{Z}_{2}$
shares the same identity component as the ordinary Lorentz group $O(3,1)$ but has disconnected components generated by $\Lambda_{\infty}(\theta,\phi)$ and $\Lambda_{-\infty}(\theta,\phi)$ in place of parity and time-reversal. When translations are included, this leads to an extended Poincaré group $\mathcal{P}_{\mathrm{ext}}=\mathbb{R}^{1,3}\rtimes\mathcal{L}_{\mathrm{ext}}$ whose unitary irreducible representations (UIRs) differ from the standard Wigner classification [2] in two essential ways: (i) timelike and spacelike orbits are merged into a single tachyon/massive multiplet, and (ii) in the massless sector the stability subgroup is enlarged from Wigner’s $ISO(2)$ to
$ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}\mathbb{Z}_{2},\qquad\mathbb{Z}_{2}=\{1,-1\},\quad\Lambda_{-\infty}^{2}=I.$
The presence of this extra $\mathbb{Z}_{2}$ factor in the stabiliser of a lightlike orbit of the extended group has a direct and unavoidable representation–theoretic consequence. Every massless UIR of $\mathcal{P}_{\mathrm{ext}}$ is no longer a single forward Wigner representation, but comes as a two components object
$\pi_{\mathrm{ml}}^{\mathrm{ext}}\;\simeq\;\pi^{\text{fwd}}_{\varepsilon}\oplus\pi^{\text{bwd}}_{\varepsilon},\qquad\varepsilon=\pm 1,$
where $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ are forward and backward lightlike Wigner’s UIRs of ordinary Poincaré group, and $\varepsilon$ is the eigenvalue of the unitary operator $U(\Lambda_{\infty})$. Note that $\Lambda_{-\infty}$ is the element in the little group that fixes $p_{0}$ enlarging the stability group of lightlike orbit from $ISO(2)$ to $ISO(2)\rtimes_{\text{Ad}_{\Lambda_{-\infty}}}$, while $\Lambda_{\infty}$ exchanges the forward and backward sectors; on $\mathcal{H}_{\text{fwd}}\oplus\mathcal{H}_{\text{bwd}}$ the latter is represented by an off-diagonal involution whose eigenvectors have eigenvalues $\varepsilon=\pm 1$. In consequence of this the massless sector always carries an additional binary internal label $\varepsilon=\pm 1$ attached to the pair $(\pi^{+},\pi^{-})$.
In section 2.3 we show that this binary structure realises, in a spacetime–geometric way, the algebraic pattern usually associated with entangled qubit pairs. The representation space of $\pi_{\mathrm{ml}}^{\mathrm{ext}}$ can be written as a direct sum
$\mathcal{H}_{\oplus}=\mathcal{H}^{\text{fwd}}\oplus\mathcal{H}^{\text{bwd}},$
where the above summands carry forward and backward massless Wigner UIRs respectively, while the action of $U(\Lambda_{\infty})$ exchanges the two sectors up to a sign determined by $\varepsilon$. In particular we show that there exist:
  • a unitary sector isometry
[…]
    where $\mathcal{H}$ is isomorphic to $\mathcal{H}^{\text{fwd}}\simeq\mathcal{H}^{\text{bwd}}$, and
  • a natural *–homomorphism of observable algebras
    $\iota:\ \mathcal{B}(\mathcal{H})\otimes M_{2}(\mathbb{C})\ \longrightarrow\ \mathcal{B}(\mathcal{H}_{\oplus}),$
such that for every state $\Psi\in\mathcal{H}_{\oplus}$ and for every local observable $A\otimes B$ one has
$\langle\Psi,\ \iota(A\otimes B)\,\Psi\rangle_{\mathcal{H}_{\oplus}}=\langle V\Psi,\ (A\otimes B)\,V\Psi\rangle_{\mathcal{H}\otimes\mathbb{C}^{2}}.$
In other words, a massless UIR of the extended Poincaré group is operationally indistinguishable, on all observables of the form $A\otimes B$, from an entangled state of two qubits where one of the qubits encodes the binary degree of freedom $\varepsilon$ associated with $U(\Lambda_{\infty})$. Entanglement in this setting is therefore not an additional structure put in by hand, but a consequence of the geometry of the extended Lorentz symmetry.
In section 3 we show that these massless representations have already been realized in laboratory exploiting entanglement in degrees of freedom of a single photon (see e.g. [4]).
We then explore these photonic states as a resource for quantum computation. Quantum computation is usually presented by postulating a two–level system as a primitive. In photonics this is often realised by selecting a two–dimensional code subspace inside a larger physical Hilbert space (e.g. polarisation within a multi–mode field, dual–rail in spatial modes, time bins, etc.). In section 4 we show that the representation space of quantum states describing the entanglement between degrees of freedom of a single photon (e.g. helicity and direction of propagation) already identify the quantum states of a logical qubit. Operationally, the resulting qubit is implemented as a single photon entanglement (SPE) state between propagation direction and polarisation and the two logical basis states appear as Bell–like states in a fixed two–mode single–photon manifold. The qubit code space $\mathcal{C}$ is operatively constructed by providing the single photon entanglement logical basis with two physically standard one–parameter optical controls (a relative phase shifter and a coherent two–mode coupler) that, when represented in the logical basis, generate all single–qubit unitaries $SU(2)$ on the code space $\mathcal{C}$. This choice immediately provides a canonical Bloch sphere of pure logical states and a CPTP instrument on the tensor product space of two of such Bloch spheres whose Choi state provides an entangling resource for universal quantum computation. In particular, in section 5, we will introduce a two–photons polarisation parity measurement and show that expressed in logical basis is equivalent to a Pauli–product measurement on the tensor product of two single photon entanglement (SPE) qubits $\mathcal{C}_{A}\otimes\mathcal{C}_{B}$, from which an entangling logical gate follows via a standard gate–teleportation construction. From this it follows universality of the computational model.
2 Massless Representations of Extended Poincarè Group
2.1 Unitary Irreducible Representations (UIRs) Classification
The construction in [1] begins by considering superluminal boosts of velocity $v>c$ and taking the limit $|v|\to\infty$ while keeping the direction fixed. This yields an involutive Lorentz transformation $\Lambda_{\infty}(\theta,\phi)$ whose action in the $(t,\vec{x})$ coordinates is
$t\;\mapsto\;\hat{n}\cdot\vec{x},\qquad\hat{n}\cdot\vec{x}\;\mapsto\;t,\qquad\hat{n}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta).$
Conjugation by $\Lambda_{\infty}$ is a nontrivial automorphism of the proper orthochronous group ${SO}(3,1)^{+}_{\uparrow}$ and generates the extension
$\mathcal{L}_{\mathrm{ext}}\cong\ ({SO}(3,1)^{+}_{\uparrow}\rtimes_{{\text{Ad}}_{\Lambda_{\infty}}}\mathbb{Z}_{2})\times\mathbb{Z}_{2}$
where $\mathbb{Z}_{2}$ is $\{$-1,1}\}, the nontrivial automorphism $\text{Ad}_{\Lambda_{\infty}}$ is conjugation by $\Lambda_{\infty}(\theta,\phi)$ and conjugation with matrices corresponding to different values $(\theta^{\prime},\phi^{\prime})$ produces equivalent group extensions. Including translations we defined the extended Poincaré group
$\mathcal{P}_{\text{ext}}\;=\;\mathcal{T}\rtimes\mathcal{L}_{\text{ext}}$ (1)
with multiplication
$(h,a)\cdot(h^{\prime},a^{\prime})\;=\;\bigl(hh^{\prime},\,a+ha^{\prime}\bigr),$
where $\mathcal{T}\cong\mathbb{R}^{4}$ is translation group, $h\in\mathcal{L}_{\text{ext}}$ acts on $\mathcal{T}$ by the standard linear action on $\mathbb{R}^{4}$.
Exploiting the fact that
$\mathcal{P}_{\text{ext}}=((\mathcal{T}\rtimes{SO}(3,1)^{+}_{\uparrow})\rtimes_{{\text{Ad}}_{\Lambda_{\infty}}}\mathbb{Z}_{2})\times\mathbb{Z}_{2}=(\mathcal{P}_{0}\rtimes_{{\text{Ad}}_{\Lambda_{\infty}}}\mathbb{Z}_{2})\times\mathbb{Z}_{2}$ (2)
we are able to classify unitary irreducible representations (UIRs) of the extended group $\mathcal{P}_{\text{ext}}$ from the induced action of the extension group:
$Z=\{I,-I,\Lambda_{\infty},-\Lambda_{\infty}\}$ (3)
on UIRs of Poincaré group $\mathcal{P}_{0}$. In order to explain this we call $N\equiv\mathcal{P}_{0}=(\mathcal{T}\rtimes SO(3,1)^{+}_{\uparrow})$ and define
N^={Wigner’s UIRs of $\hat{N}$$\{\text${Wigner's UIRs of N}\} (4)
An element $z\in Z$ acts on $N$ via automorphisms, namely its action on a given element of the Poincaré group $n=(h,a)\in N$ with $h\in SO(3,1)^{+}_{\uparrow}$:
$n^{\prime}=z\cdot n$ (5)
is such that it exists an automorphism of $N$, $\alpha_{z}(n)$ such that
$n^{\prime}=\alpha_{z}(n)$ (6)
where $\alpha_{z}$ is defined as:
$\alpha_{z}(h,a)=\big(zhz^{-1},\,z\!\cdot\!a\big),\qquad z\in Z$ (7)
This induces an action on the set of characters (i.e. UIRs) of $\mathcal{P}_{0}$:
$z\cdot\pi(n)=\pi(\alpha_{z}(n))\;\;\;\forall n\in N$ (8)
The set $\hat{N}$ is constituted by UIRs of Poincaré group which have been classified by Wigner in [2]. UIRs of Poincaré group are classified by the value of the four momentum vector $p=p_{\mu}$ together with the value of the invariant lenght $|L_{s}\cdot p|^{2}$ associated to the orbit
$O_{p}=\{L_{s}\cdot p|L_{s}\in SO(3,1)^{+}_{\uparrow}\}$ (9)
$p$ is called the representative of the induced UIR of Poincaré group. Given choice of $p_{\mu}$ an $SO(3,1)^{+}_{\uparrow}$-orbit representative in Wigner’s classification, its invariant could be $p^{\mu}p_{\mu}\neq 0$ or $p^{\mu}p_{\mu}=0$. Choosing a non massless ($p^{\mu}p_{\mu}\neq 0$) representative $p_{\mu}$ we have:
  • Massive Representations: Unitary irreps of SO(3) for time-like orbits
    $\pi^{+}_{\text{mass}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}>0\;\;\;p_{0}>0\}$ (10)
    $\pi^{-}_{\text{mass}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}>0\;\;\;p_{0}<0\}$ (11)
  • Tachyonic Representations: Unitary irreps of SO(2,1) for space-like orbits
    $\pi_{\text{tach}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}<0\;\;\;\}$ (12)
Choosing a massless ($p^{\mu}p_{\mu}=0$) we have
  • Massless Representations: Unitary irreps of ISO(2) for light-like orbits
    $\pi_{\text{massless}}^{\text{fwd}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}=0\;\;p^{0}>0\}$ (13)
    $\pi_{\text{massless}}^{\text{bwd}}:=\text{UIRs of }\;\;\{\ L\in SO(3,1)^{+}_{\uparrow}|L^{\mu}_{\nu}p_{\mu}=p_{\nu}\;\;\;p^{\mu}p_{\mu}=0\;\;p^{0}<0\}$ (14)
Now, with reference to the action in (8) define, given a representative $\pi\in\hat{N}$, its $Z$-orbit:
$O_{\pi}=\{z\pi\;\;|\;\;z\in Z\}$ (15)
and the stabilizers
$Z_{\pi}=\{z\in Z\;\;|\;\;z\pi(n)=\pi(n)\}$ (16)
where $n$ is in $N\equiv\mathcal{P}_{0}$.
We have two different possibilities [3].
If $Z_{\pi}=\{e\}\;\;\forall\pi(n)\;\in\;\hat{N}$ with $e$ identity element of $\pi$ then $\exists\;U(z)$ such that
$U(z)\pi(n)U(z)=\pi(z^{-1}nz)\;\;\;\forall z\in Z\;\;$ (17)
and the UIRs of $\mathcal{P}_{0}$ (namely the set $\hat{N}$) consist of UIRs of the extended group too. In this case UIRs which are inequivalent as representations of $\mathcal{P}_{0}$ become equivalent for $\mathcal{P}_{\text{ext}}$. This is due to the fact that orbits of $SO(3,1)^{+}_{\uparrow}$ in momentum space preserve $p^{\mu}p_{\mu}$ with its sign while in $\mathcal{L}_{\text{ext}}$ the sign can change.
If, on the contrary, $Z_{\pi}\neq\{e\}$ for some $\pi$ then its elements may give rise to inequivalent UIRs depending on the action of z $\in Z_{\pi}$ on $\pi\in\hat{N}$.
2.2 Massless UIRs
Derivation of massless UIRs of the extended group are detailed in [1]. Here we summarize the relevant conceptual pillars. It is considered the lightlike zz–orbit
$\mathcal{O}_{0}\;=\;\{\,Lp_{0}\;|\;L\in SO(3,1)^{+}_{\uparrow},\ p_{0}^{2}=0\,\},$ (18)
with the standard representative chosen as
$p_{0}=(\omega,\omega\sin\theta\cos\phi,\omega\sin\theta\sin\phi,\omega\cos\theta),\qquad\omega>0.$ (19)
The usual Wigner classification for the ordinary Poincaré group $\mathcal{P}_{0}$ yields two massless UIRs, denoted
$\pi^{\text{fwd}},\ \pi^{\text{bwd}},$
corresponding to forward and backward light–cones ($p_{0}^{0}>0$ and $p_{0}^{0}<0$). They are both induced from unitary irreps of the Euclidean group $ISO(2)$ (helicity or continuous–spin class) and are inequivalent as representations of $\mathcal{P}_{0}$.
In the extended Lorentz group $\mathcal{L}_{\mathrm{ext}}$ the matrix $\Lambda_{-\infty}=-\Lambda_{\infty}$ satisfies $\Lambda_{-\infty}p_{0}=p_{0},$ hence it lies in the stability subgroup of the chosen representative. Consequently, the geometric little group at $p_{0}$ in $\mathcal{L}_{\mathrm{ext}}$ is
$ISO(2)\rtimes\{I,\Lambda_{-\infty}\}.$ (20)
Let $U_{0}$ be a unitary representation of the little group $ISO(2)$ associated with either helicity or continuous–spin. The induced Poincaré action on wavefunctions $\psi:\mathcal{O}_{0}\to\mathcal{H}_{0}$ has the usual form
$\bar{U}(a,h)\psi\,](p)\;=\;e^{\,i\,p\cdot a}\;U_{0}\!\big(s_{0}(h,p)\big)\;\psi(h^{-1}p),\qquad(a,h)\in\mathcal{P}_{0},$ (21)
and this construction gives the standard massless Wigner UIRs $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ depending on the choice of representative on the forward or backward orbit.
In contrast to $\Lambda_{-\infty}$, the transformation $\Lambda_{\infty}$ maps $p_{0}$ to the backward representative
$\Lambda_{\infty}(\theta,\phi)\,p_{0}=-p_{0},$ (22)
thus interchanging the forward and backward light–cones. At the level of massless UIRs of $\mathcal{P}_{\text{ext}}$, $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ are equivalent since the unitary operator representing $U(\Lambda_{\infty})$ is an intertwiner between $\pi_{\text{fwd}}$ and $\pi_{\text{bwd}}$
$U(\Lambda_{\infty})\;\pi^{\text{fwd}}(n)\;U(\Lambda_{\infty})^{-1}\;=\;\pi^{\text{fwd}}\bigl(\Lambda_{\infty}^{-1}n\Lambda_{\infty}\bigr)\;=\;\pi^{\text{bwd}}(n),\qquad\forall n\in\mathcal{P}_{0}.$ (23)
Thus, under the action of the discrete factor $Z_{(\theta,\phi)}=\{I,-I,\Lambda_{\infty},-\Lambda_{\infty}\}$ on the dual space $\widehat{\mathcal{P}_{0}}$, (i.e. the set of Wigner’s UIRs) the set $\Pi\equiv\{\pi^{\text{fwd}},\pi^{\text{bwd}}\}$ constitute a single equivalence class of UIRs in which the transformation $\Lambda_{\infty}$ is a non trivial stabilizer, namely:
$\Lambda_{\infty}\Pi=\Pi$ (24)
As a consequence, the induced massless UIRs of the extended Poincaré group $\mathcal{P}_{\mathrm{ext}}$ do not live on a single copy of $\mathcal{H}_{0}$, but rather on the direct sum
$\mathcal{H}_{\oplus}\;=\;\mathcal{H}_{\text{fwd}}\oplus\mathcal{H}_{\text{bwd}},$ (25)
where $\mathcal{H}_{\text{fwd}}$ and $\mathcal{H}_{\text{bwd}}$ carry $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$, respectively. Writing $\Psi=(\psi_{\text{fwd}},\psi_{\text{bwd}})\in\mathcal{H}_{\oplus}$, the extended action of $(a,h)\in\mathcal{P}_{0}$ is block–diagonal:
$\bar{U}(a,h)\Psi\,](p)=\bigl([\bar{U}_{\text{fwd}}(a,h)\psi_{\text{fwd}}](p),[\bar{U}_{\text{bwd}}(a,h)\psi_{\text{bwd}}](p)\bigr),$ (26)
with each component governed by (21).
The discrete elements of $\mathcal{L}_{\mathrm{ext}}$ act as follows. For $\Lambda_{\infty}$ we choose an operator $\bar{U}(\Lambda_{\infty})$ on $\mathcal{H}_{\oplus}$ of the form
$\bar{U}(\Lambda_{\infty})\Psi\,](p):=\bigl(C\,\psi_{\text{bwd}}(\Lambda_{\infty}^{-1}p),\;C^{-1}\,\psi_{\text{fwd}}(\Lambda_{\infty}^{-1}p)\bigr),$ (27)
where $C$ implements the intertwining between $\pi^{\text{fwd}}$ and $\pi^{\text{bwd}}$ as above and its representation depends on the massless representation choosen for Wigner’s little group $ISO(2)$. Throughout this paper we will use helicity representation as working hypothesis and set $C=I_{\mathcal{H}}$ choosing the identification $\mathcal{H}\simeq\mathcal{H}_{\text{bwd}}\simeq\ H_{\text{fwd}}$. Using a section $k_{0}$ adapted as
$k_{0}(\Lambda_{\infty}^{-1}p)\;=\;\Lambda_{\infty}^{-1}\,k_{0}(p),$ (28)
and the covariance relation
$s_{0}\!\big(\Lambda_{\infty}h\Lambda_{\infty}^{-1},\,p\big)\;=\;\Lambda_{\infty}\,s_{0}\!\big(h,\Lambda_{\infty}^{-1}p\big)\,\Lambda_{\infty}^{-1},$ (29)
we have:
$\bar{U}(\Lambda_{\infty})\,\bar{U}(a,h)\,\bar{U}(\Lambda_{\infty})^{-1}\;=\;\bar{U}\!\big(\Lambda_{\infty}a,\ \Lambda_{\infty}h\Lambda_{\infty}^{-1}\big),$ (30)
thus $\bar{U}(\Lambda_{\infty})$ correctly represents the extension.
Similarly, for I-I we may choose
$\bar{U}(-I)\Psi\,](p):=\bigl(\psi_{\text{bwd}}(-p),\,\psi_{\text{fwd}}(-p)\bigr),$ (31)
which exchanges forward and backward sectors and satisfies
$\bar{U}(-I)\,\bar{U}(a,h)\,\bar{U}(-I)^{-1}\;=\;\bar{U}\!\big(-a,\ (-I)h(-I)^{-1}\big).$ (32)
The only further constraint on $\bar{U}(\Lambda_{\infty})$ is that $\bar{U}(\Lambda_{\infty})^{2}=\mathbb{I}.$ This restricts $C$ in (27) to a unitary with $C^{2}=\mathbb{I}$ on $\mathcal{H}_{0}$, so that its eigenvalues are $\pm 1$. Different signs for the eigenvalues of $C$ lead to two inequivalent massless UIRs of the extended group for each standard Wigner class (helicity or continuous–spin). In this sense, the lightlike orbit supports a doublet representation
$\pi^{\mathrm{ext}}_{\varepsilon}\;\sim\;\pi^{\text{fwd}}_{\varepsilon}\ \oplus\ \pi^{\text{bwd}}_{\varepsilon},\qquad\varepsilon=\pm 1,$
where $\varepsilon$ labels the choice of sign in the representation of the extended little group of lightlike momenta representatives.
2.3 Operational Equivalence of Massless UIRs with Entangled Two–Qubits States
The two states:
$\Psi_{\varepsilon=\pm 1}=\frac{1}{\sqrt{2}}(\psi_{\text{fwd}}\oplus\pm\psi_{\text{bwd}})$ (33)
are eigenstates of $\bar{U}(\Lambda_{\infty})$ corresponding to eigenvalues $\pm 1$. We now choose an isomorphism $\mathcal{H}\simeq\mathcal{H}_{\text{fwd}}\simeq\mathcal{H}_{\text{bwd}}$, a set of orthonormal vectors $|\text{fwd}\rangle,|\text{bwd}\rangle$ spanning a two–dimensional complex space $\mathcal{K}\simeq\mathbb{C}^{2}$ and define the sector isometry in the represenation space (25):
$V:\mathcal{H}_{\oplus}\to\mathcal{H}\otimes\mathcal{K},\qquad V(\Psi_{\varepsilon=\pm 1})=\psi_{\text{fwd}}\otimes|\text{fwd}\rangle\pm\psi_{\text{bwd}}\otimes|\text{bwd}\rangle.$ (34)
This is unitary and preserves inner products.
Let $\mathcal{B}(\mathcal{H})$ be the algebra of bounded operators on $\mathcal{H}$ and $\mathcal{B}(\mathcal{K})$ the algebra of all $2\times 2$ matrices acting on the internal qubit $\text{span}\{|\text{fwd}\rangle,|\text{bwd}\rangle\}$. For any $A\in\mathcal{B}(\mathcal{H})$ and $B=[b_{ij}]\in\mathcal{B}(\mathcal{K})$ define the unital *–homomorphism
$\iota(A\otimes B)=\begin{pmatrix}b_{00}A&b_{01}A\\ b_{10}A&b_{11}A\end{pmatrix},\qquad\iota:\mathcal{B}(\mathcal{H})\otimes B(\mathcal{K})\;\to\;\mathcal{B}(\mathcal{H}_{\oplus}).$ (35)
It preserves products, adjoints, and the identity operator.
A direct calculation shows that for any linear superposition $\Psi$ of states in (33) we have
$\big\langle\Psi,\,\iota(A\otimes B)\,\Psi\big\rangle=\big\langle V\Psi,\,(A\otimes B)\,V\Psi\big\rangle.$ (36)
Therefore all observable predictions of the $\mathcal{P}_{\mathrm{ext}}$ massless UIR coincide with those of an entangled state of two qubits. The states in (33) correspond under $V$ to Bell-like states.
Thus entanglement is not added by hand but emerges from the geometry of the extended Lorentz symmetry.
In block matrix form (with respect to the decomposition $\mathcal{H}\oplus\mathcal{H}$) the action of $U(\Lambda_{\infty})$ in (27) is
$U(\Lambda_{\infty})=\begin{pmatrix}0&I_{\mathcal{H}}\\ I_{\mathcal{H}}&0\end{pmatrix}$, (37)
On the tensor product space we consider the action of $I\otimes\sigma_{x}$ where $\sigma_{x}$ is the Pauli matrix acting on $\mathcal{K}\equiv\text{span}\{|\text{fwd}\rangle,|\text{bwd}\rangle\}$ as:
$\sigma_{x}|\text{fwd}\rangle=|\text{bwd}\rangle,\quad\sigma_{x}|\text{bwd}\rangle=|\text{fwd}\rangle.$
and thus, in this basis, has the form:
$\sigma_{x}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, (38)
We now compare the action of $I\otimes\sigma_{x}$ on $V(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})$ with the action of $V$ on $U(\Lambda_{\infty})(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})$.
The first is:
$(I\otimes\sigma_{x})\,V(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})=(I\otimes\sigma_{x})\big(\psi_{\text{fwd}}\otimes|\text{fwd}\rangle+\psi_{\text{bwd}}\otimes|\text{bwd}\rangle\big)$ (39)
On the other hand,
$U(\Lambda_{\infty})(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})=\psi_{\text{bwd}}\oplus\psi_{\text{fwd}},$ (40)
so
$V\big(U(\Lambda_{\infty})(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})\big)=V(\psi_{\text{bwd}}\oplus\psi_{\text{fwd}})=\psi_{\text{bwd}}\otimes|\text{fwd}\rangle+\psi_{\text{fwd}}\otimes|\text{bwd}\rangle.$ (41)
Therefore
$(I\otimes\sigma_{x})\,V(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})\;=\;V\big(U(\Lambda_{\infty})(\psi_{\text{fwd}}\oplus\psi_{\text{bwd}})\big)$ (42)
Since this must hold for all wavefunctions in $\mathcal{H}_{\oplus}$ we conclude that, via the isometric identification $V$, the sector–swap operator $U(\Lambda_{\infty})$ on $\mathcal{H}\oplus\mathcal{H}$ is unitarily equivalent to the operator $I\otimes\sigma_{x}$ on $\mathcal{H}\otimes\mathbb{C}^{2}$.
3 Experimental Access to Massless UIRs of Extended Poincarè Group
In this section we clarify the operational meaning of the additional binary degree of freedom $\varepsilon=\pm 1$ associated with $U(\Lambda_{\infty})$, and we outline a concrete single–photon interferometric implementation in standard quantum optics. The goal is to connect the abstract decomposition
$\mathcal{H}_{\mathrm{ml}}^{\mathrm{ext}}\simeq\mathcal{H}_{\text{fwd}}\oplus\mathcal{H}_{\text{bwd}}$
with experimentally accessible photon modes.
3.1 Specialisation to a Fixed Direction and Single–Mode Photons
In the extended Poincaré picture the extra structure in the massless sector comes from the choice of a spatial direction
$\hat{n}(\theta,\phi)=(\sin\theta\cos\phi,\ \sin\theta\sin\phi,\ \cos\theta),$
[…]
$p_{+}=\omega(1,0,0,1),\qquad p_{-}=\omega(1,0,0,-1),$
which are the forward and backward representatives associated with this choice of $\hat{n}$. Rotating the entire setup in space would simply correspond to choosing a different pair $(\theta,\phi)$ in the abstract construction. So $p_{0}=\omega(1,0,0,1),$ with $\Lambda_{\infty}=\Lambda_{\infty}(0,0)$ an involution exchanging the time coordinate $t$ with the spatial coordinate $z$. we can represent a fixed helicity component by kets
$|p_{+}\rangle\in\mathcal{H}_{\text{fwd}},\qquad|p_{-}\rangle\in\mathcal{H}_{\text{bwd}},$ (43)
corresponding to the two lightlike directions along +z+z and z-z. Under the isomorphism $\mathcal{H}_{\text{fwd}}\simeq\mathcal{H}_{\text{bwd}}\simeq\mathcal{H}$, these are mapped to the “direction basis”
$|+\rangle,\ |$-$\rangle\in\mathcal{H}$
Once the spatial direction $\hat{n}$ is fixed, the remaining geometric degree of freedom in the theory is the eigenvalue $\varepsilon=\pm 1$ of $U(\Lambda_{\infty})$ acting on the two–component massless UIR. In the optical realisation this binary label is encoded in the relative phase between the forward and backward components of the photon state.
Restricting to a single momentum mode and a fixed helicity, and using the identifications
[…]
$|\text{fwd}\rangle\equiv|H\rangle,\qquad|\text{bwd}\rangle\equiv|V\rangle,$
the prepared single–photon state at the output of the interferometer (with a controllable phase shift in one arm) can be written as
$|\Psi(\varphi)\rangle=\frac{1}{\sqrt{2}}\Big(|+\rangle\otimes|H\rangle+e^{i\varphi}\,|-\rangle\otimes|V\rangle\Big),$ (44)
where $\varphi$ is the experimentally tunable interferometric phase (e.g. adjusted with a phase shifter or by varying the optical path length in one arm).
In the abstract representation–theoretic description, the $U(\Lambda_{\infty})$–eigenstates are precisely the symmetric and antisymmetric superpositions
$|\Psi_{\varepsilon=\pm 1}\rangle=\frac{1}{\sqrt{2}}\Big(|+\rangle\otimes|H\rangle\pm|-\rangle\otimes|V\rangle\Big),$ (45)
which are obtained from (44) by setting
$\varphi=\begin{cases}0\mod 2\pi,&\varepsilon=+1,\\[3.0pt] \pi\mod 2\pi,&\varepsilon=-1.\end{cases}$ (46)
Thus the theory predicts that controlling the eigenvalue sector $\varepsilon$ is operationally equivalent to controlling the interferometric phase $\varphi$ between the two counterpropagating components.
3.2 Preparation Stage
A preparation scheme consistent with the above identification is:
  1. 1.
    Use a heralded single–photon source to prepare photons in a well-defined wave packet peaked around the four–momentum $p_{0}=\omega(1,0,0,1)$, with fixed helicity (e.g. right–handed circular polarisation).
  2. 2.
    Send the photon into an interferometer (e.g. a Michelson or Mach–Zehnder configuration) aligned along the zz–axis, with a highly reflecting mirror in one arm so that the output state contains a coherent superposition of forward and backward propagation along $\pm\hat{z}$. At the level of the abstract representation, this realises a superposition of the two sector basis vectors $|+\rangle$ and $|$-$\rangle$.
  3. 3.
    Insert polarisation optics in the two arms so that the component travelling in the $\hat{z}$ direction emerges with polarisation $|H\rangle$ and the component travelling in the $\hat{z}$ direction emerges with polarisation $|V\rangle$. Relative phases between the two paths can be tuned with phase shifters to realise
    $\frac{1}{\sqrt{2}}\bigl(|+\rangle\otimes|H\rangle+\varepsilon\,|-\rangle\otimes|V\rangle\bigr),$
    with $\varepsilon=\pm 1$ determined by the interferometric phase (see 46).
From the viewpoint of the extended Poincaré group, the two classes of preparations with relative phase $\varepsilon=\pm 1$ are interpreted as populating the two eigenvalue sectors of $U(\Lambda_{\infty})$ associated with the lightlike orbit in the direction $\hat{n}=(0,0,1)$.
3.3 Measurement Stage and Dependence on $\varepsilon$
To detect $\varepsilon$ we must measure an observable whose effective action, in the abstract description, is $\iota(A\otimes\sigma_{x})$ (see (35)) with an $A$ that mixes the forward/backward sectors. A natural choice is the Pauli operator
$A=\sigma_{x}^{(\mathrm{dir})}=|+\rangle\langle-|+|-\rangle\langle+|$
on the direction subspace. On the polarization subspace (encoding the basis $|\text{fwd}\rangle,|\text{bwd}\rangle$) we are forced to measure $\sigma_{x}^{(\mathrm{pol})}$. In the two–qubit picture the corresponding correlation observable is
$O_{XX}=\sigma_{x}^{(\mathrm{dir})}\otimes\sigma_{x}^{(\mathrm{pol})}.$
A straightforward calculation in the basis
$\{|+\rangle\otimes|H\rangle,\ |+\rangle\otimes|V\rangle,\ |-\rangle\otimes|H\rangle,\ |-\rangle\otimes|V\rangle\}$
yields, for the state (45),
$\langle\Psi_{\varepsilon}|\,O_{XX}\,|\Psi_{\varepsilon}\rangle=\varepsilon.$ (47)
Thus the sign of the correlation between (i) direction measured in the XX–basis of the forward/backward sectors and (ii) polarisation measured in the XX–basis of the internal qubit is exactly the eigenvalue $\varepsilon$ of $U(\Lambda_{\infty})$.
In the optical realisation, $O_{XX}$ can be implemented with standard components:
  • A 50:50 beam splitter acting on the counterpropagating modes along $\pm\hat{z}$ interferes $|+\rangle$ and $|$-$\rangle$ into two output ports corresponding to the eigenstates of $\sigma_{x}^{(\mathrm{dir})}$.
  • In each output port a half–wave plate at $45^{\circ}$ followed by a polarising beam splitter measures polarisation in the $\{|+\rangle_{X},|-\rangle_{X}\}$ basis (eigenbasis of $\sigma_{x}^{(\mathrm{pol})}$).
  • Single–photon detectors at the exits record joint outcomes $(x_{\mathrm{dir}},x_{\mathrm{pol}})\in\{\pm 1\}\times\{\pm 1\}$, from which one reconstructs the correlation
    $E_{XX}=\sum_{x_{\mathrm{dir}},x_{\mathrm{pol}}=\pm 1}x_{\mathrm{dir}}x_{\mathrm{pol}}\,P(x_{\mathrm{dir}},x_{\mathrm{pol}}),$
    which in the idealised extended theory reproduces (47).
3.4 Interpretation and Falsifiability
From the viewpoint of standard quantum optics, the above is a textbook single–photon entanglement experiment between a path–like degree of freedom (here encoding the two sectors associated with the lightlike direction $\hat{n}$) and polarisation. What is nontrivial in the present work is the representation–theoretic interpretation: the same setup constitutes a local tomography of the binary degree of freedom $\varepsilon$ associated with $U(\Lambda_{\infty})$ in the massless UIR of $\mathcal{P}_{\mathrm{ext}}$.
The extended theory predicts that:
  • There exist preparations and measurements, compatible with the extended symmetry, in which the sign of the correlation $E_{XX}$ can be associated with the eigenvalue $\varepsilon$ of $U(\Lambda_{\infty})$ for a fixed lightlike direction $\hat{n}$, and in which switching sector corresponds to a flip $E_{XX}\to-E_{XX}$ without changing local marginals on $\mathcal{H}_{+}$ or $\mathcal{H}_{-}$.
  • If, in all such optical realisations, only one effective value of $\varepsilon$ is observed (e.g. $E_{XX}$ never changes sign in regimes where the extended theory predicts that both sectors should be accessible), then the massless sector of the extended Poincaré group would be empirically constrained or falsified.
In practice, imperfections such as losses, detector inefficiency and mode mismatch will reduce $|E_{XX}|$ below unity; one would then look for a statistically significant change of sign of the reconstructed correlation as a function of controlled preparation parameters. Nevertheless, the conceptual link is clear: by fixing the direction $\hat{n}$ of $\Lambda_{\infty}$, identifying the forward/backward sectors with counterpropagating photon modes along $\pm\hat{n}$, and performing local tomography on the resulting two–qubit system, one can in principle probe the geometric degree of freedom $\varepsilon$ and thus test the proposed geometric origin of entanglement for massless fields.
3.5 Relation to Existing Single–Particle Entanglement Experiments
The state and measurement we propose in this section are, from the point of view of standard quantum optics, instances of single–particle entanglement between two internal degrees of freedom. In particular, the Bell–like state
$|\Psi_{\varepsilon}\rangle\;\simeq\;\frac{1}{\sqrt{2}}\bigl(|+\rangle\otimes|H\rangle+\varepsilon\,|-\rangle\otimes|V\rangle\bigr),\qquad\varepsilon=\pm 1,$ (48)
is a special case of the generic path–polarisation entangled state of a single photon, where $|+\rangle,|$-$\rangle$ encode two spatial modes and $|H\rangle,|V\rangle$ encode two polarisation modes. Correlations of Pauli–type observables such as $\sigma_{x}^{(\mathrm{dir})}\otimes\sigma_{x}^{(\mathrm{pol})}$ have been measured in several experiments.
For example, Fiorentino et al. implement a deterministic controlled–NOT gate acting on two qubits carried by a single photon: one qubit is encoded in polarisation, the other in the spatial mode (momentum) of the photon [4]. Their preparation and analysis stages generate and characterise states of the form
$|\Phi(\phi)\rangle=\frac{1}{\sqrt{2}}\bigl(|0\rangle_{\mathrm{path}}\otimes|H\rangle+e^{i\phi}\,|1\rangle_{\mathrm{path}}\otimes|V\rangle\bigr),$ (49)
which are mathematically equivalent to our $|\Psi_{\varepsilon}\rangle$ for phases $\phi=0,\pi$ and with an appropriate identification of $|0\rangle_{\mathrm{path}},|1\rangle_{\mathrm{path}}$ with $|+\rangle,|$-$\rangle$. Full two–qubit tomography in mutually unbiased bases (including the $X$ basis on each qubit) is performed, so that correlations proportional to $\langle\sigma_{x}^{(\mathrm{path})}\otimes\sigma_{x}^{(\mathrm{pol})}\rangle$ are effectively accessed [4].
Similarly, Bera et al. propose a protocol in which intra–photon entanglement between path and polarisation is swapped to inter–photon entanglement using linear optics [5]. Their initial single–photon resource is again a maximally entangled state between path and polarisation, structurally identical to $|\Psi_{\varepsilon}\rangle$, and the protocol requires the ability to prepare, transform and analyse such states in various local bases.
[…]
From the purely operational point of view, these experiments show that
  • a state preparation equivalent (up to local unitaries) to $|\Psi_{\varepsilon}\rangle$ is experimentally standard,
  • joint measurements of Pauli observables on the two internal degrees of freedom—including $X\otimes X$–type correlations—are feasible and have been performed, and
  • the observed correlations agree with the quantum predictions for a maximally entangled two–qubit state within experimental accuracy.
The novelty of the present work does not lie in the optical technology required to prepare and measure such single–photon two–qubit states, but in the representation–theoretic interpretation. In the above experiments, the two qubits are treated as abstract, independent Hilbert–space degrees of freedom (path, polarisation, spin), and their entanglement is a kinematical feature of the chosen encoding. In our framework, by contrast,
  1. 1.
    the “path” degree of freedom $|+\rangle,|$-$\rangle$ is identified with the forward/backward lightlike sectors associated with a fixed null momentum orbit of the extended Poincaré group, and
  2. 2.
    the second qubit corresponds to the binary internal label $\varepsilon=\pm 1$ associated with the eigenvalues of the superluminal involution $U(\Lambda_{\infty})$ in the massless UIR of $\mathcal{P}_{\mathrm{ext}}$.
In other words, existing experiments already implement the same mathematical structure—a two–qubit Bell state and local Pauli measurements—but they do not test the geometric claim that this structure arises from the extended Lorentz symmetry and that the internal qubit is nothing but the eigenvalue sector of $U(\Lambda_{\infty})$. Our proposed experiment is therefore best viewed as a reinterpretation and a targeted adaptation of standard single–photon entanglement setups, designed to perform a tomography of the $\varepsilon$ degree of freedom tied to the extended Poincaré representation, rather than a test of quantum mechanics itself.
4 Single Photon Entanglement Qubit
In this section we show that entangled two qubits states introduced in the previous section, together with a tunable relative phase shifter and a two modes photons coupler, can be used to operatively construct the manifold of possible states of a single qubit, i.e. a Bloch sphere. We will call the corresponding idealized system single photon entanglement qubit. In section 4.1 we will encode perfectly distinguishable states in the two eigenstates of $U(\Lambda_{\infty})$, while in section 4.2 we will show that the chosen encoding allows access of the whole Bloch sphere by means of a tunable relative phase shifter and a photonic two modes coupler.
4.1 Encoding
The parameters $(\theta,\phi)$ entering $\Lambda_{\infty}(\theta,\phi)$ have the meaning of selecting a propagation direction in the laboratory rest frame: they fix a spatial axis
$\hat{n}(\theta,\phi)=(\sin\theta\cos\phi,\ \sin\theta\sin\phi,\ \cos\theta)$
and therefore fix what we call the two counterpropagating “direction” modes $|+\rangle$ and $|$-$\rangle$ along $\pm\hat{n}$. Once the lab axis is fixed, the associated involution $\bar{U}(\Lambda_{\infty})$ has two eigenvectors who span a two dimensional Hilbert space which can be written as Bell states. In order to see this define:
$|\Phi_{+H}\rangle:=|+\rangle\otimes|H\rangle,\qquad|\Phi_{-V}\rangle:=|-\rangle\otimes|V\rangle,$
where $|+\rangle,|$-$\rangle$ label propagation direction along $\pm\hat{n}$ and $|H\rangle,|V\rangle$ label two orthogonal internal modes (e.g. polarisation). The relevant two-mode physical subspace is
$\mathcal{H}_{\mathrm{2m}}:=\mathrm{span}\{|\Phi_{+H}\rangle,|\Phi_{-V}\rangle\}\subset\mathbb{C}^{2}_{\rm dir}\otimes\mathbb{C}^{2}_{\rm pol}.$ (50)
In this basis the state in (48) becomes:
$|\Psi_{\pm}\rangle\;:=\;\frac{1}{\sqrt{2}}\Big(|\Phi_{+H}\rangle\pm|\Phi_{-V}\rangle\Big),$ (51)
We take them as logical basis vectors and define the code space
$\mathcal{C}:=\mathrm{span}\{|\Psi_{+}\rangle,|\Psi_{-}\rangle\}\cong\mathbb{C}^{2},\qquad|0_{L}\rangle:=|\Psi_{+}\rangle,\ \ |1_{L}\rangle:=|\Psi_{-}\rangle.$ (52)
Starting from the logical states defined above $|0_{L}\rangle$ and $|1_{L}\rangle$ one can define a generic state of the Bloch sphere as:
$|\psi_{L}(\vartheta,\varphi)\rangle=\cos\frac{\vartheta}{2}\,|0_{L}\rangle+e^{i\varphi}\sin\frac{\vartheta}{2}\,|1_{L}\rangle,\qquad 0\leq\vartheta\leq\pi,\ \ 0\leq\varphi<2\pi.$ (53)
In the physical two-mode embedding $\mathcal{H}_{\mathrm{2m}}=\mathrm{span}\{|\Phi_{+H}\rangle,|\Phi_{-V}\rangle\}$ this becomes
$|\psi_{L}(\vartheta,\varphi)\rangle=\frac{1}{\sqrt{2}}\Big((\cos\frac{\vartheta}{2}+e^{i\varphi}\sin\frac{\vartheta}{2})\,|\Phi_{+H}\rangle+(\cos\frac{\vartheta}{2}-e^{i\varphi}\sin\frac{\vartheta}{2})\,|\Phi_{-V}\rangle\Big),$ (54)
4.2 Physical Realization of the Bloch Sphere
We are now going to show that any state of the Bloch sphere defined above can be reached exploiting a relative phase shifter and a two mode coupler acting on the space defined in (50).
Take the definition of logical space basis:
$|0_{L}\rangle=\frac{|\Phi_{+H}\rangle+|\Phi_{-V}\rangle}{\sqrt{2}},\qquad|1_{L}\rangle=\frac{|\Phi_{+H}\rangle-|\Phi_{-V}\rangle}{\sqrt{2}}.$ (55)
This implies:
$\begin{pmatrix}|\Phi_{+H}\rangle\\ |\Phi_{-V}\rangle\end{pmatrix}=H_{L}\begin{pmatrix}|0_{L}\rangle\\ |1_{L}\rangle\end{pmatrix},\qquad H_{L}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\[2.0pt] 1&-1\end{pmatrix}$, (56)
so that the matrix of an operator $U$ in the logical basis is
$U_{L}=H_{L}^{\dagger}UH_{L}.$ (57)
Since $H_{L}^{\dagger}=H_{L}=H_{L}^{-1}$, this is just a Hadamard conjugation.
We assume two standard coherent optical controls acting on $\mathcal{H}_{2m}$:
  1. 1.
    A relative phase shifter
    $P(\alpha)=\mathrm{diag}(e^{i\alpha/2},\,e^{-i\alpha/2})=\exp\!\Big(-\frac{i}{2}\alpha\,\sigma_{z}\Big)\quad\text{in the basis }\{|\Phi_{+H}\rangle,|\Phi_{-V}\rangle\},$ (58)
    so its generator is $\frac{i}{2}\sigma_{z}$.
  2. 2.
    A two–mode coupler (directional coupler / beam splitter) mixing the two modes,
    $B(\beta)=\begin{pmatrix}\cos(\beta/2)&-\sin(\beta/2)\\ \sin(\beta/2)&\ \cos(\beta/2)\end{pmatrix}=\exp\!\Big(-\frac{i}{2}\beta\,\sigma_{y}\Big)\quad\text{in the basis }\{|\Phi_{+H}\rangle,|\Phi_{-V}\rangle\},$ (59)
    so its generator is $\frac{i}{2}\sigma_{y}$.
Using (57), the corresponding logical actions are
$P_{L}(\alpha)=H_{L}^{\dagger}P(\alpha)H_{L}=\exp\!\Big(-\frac{i}{2}\alpha\,(H_{L}^{\dagger}\sigma_{z}H_{L})\Big),\qquad B_{L}(\beta)=H_{L}^{\dagger}B(\beta)H_{L}=\exp\!\Big(-\frac{i}{2}\beta\,(H_{L}^{\dagger}\sigma_{y}H_{L})\Big).$ (60)
We now compute the Hadamard conjugations of Pauli matrices.
$H_{L}\sigma_{z}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\[2.0pt] 1&-1\end{pmatrix}\begin{pmatrix}1&0\\[2.0pt] 0&-1\end{pmatrix}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\[2.0pt] 1&1\end{pmatrix}$,
hence
$(H_{L}\sigma_{z})H_{L}=\frac{1}{2}\begin{pmatrix}1&-1\\[2.0pt] 1&1\end{pmatrix}\begin{pmatrix}1&1\\[2.0pt] 1&-1\end{pmatrix}=\frac{1}{2}\begin{pmatrix}0&2\\[2.0pt] 2&0\end{pmatrix}=\begin{pmatrix}0&1\\[2.0pt] 1&0\end{pmatrix}=\sigma_{x}.$
Similarly, one checks $H_{L}^{\dagger}\sigma_{y}H_{L}=-\sigma_{y}$ (and $H_{L}^{\dagger}\sigma_{x}H_{L}=\sigma_{z}$).
Applying Hadamard’s conjugation to $\sigma_{x}$ and $\sigma_{y}$ in (60) yields
$P_{L}(\alpha)=\exp\!\Big(-\frac{i}{2}\alpha\,\sigma_{x}\Big),\qquad B_{L}(\beta)=\exp\!\Big(-\frac{i}{2}\beta\,(-\sigma_{y})\Big)=\exp\!\Big(+\frac{i}{2}\beta\,\sigma_{y}\Big).$ (61)
Since $\beta\in\mathbb{R}$ is tunable, the sign is immaterial (replace $\beta\mapsto-\beta$). Therefore, in the logical basis, the two available one–parameter subgroups implement rotations about two distinct Bloch-sphere axes, which we identify with the logical Pauli generators:
$P_{L}(\alpha)=e^{-i\alpha X_{L}/2},\qquad B_{L}(\beta)=e^{-i\beta Y_{L}/2},\qquad(X_{L},Y_{L},Z_{L})\equiv(\sigma_{x},\sigma_{y},\sigma_{z})\ \text{on }\mathcal{C}.$ (62)
Consider the Lie algebra generated by $\{iX_{L},iY_{L}\}$. Using the Pauli commutation relations,
$X_{L},Y_{L}]=2iZ_{L},$ (63)
we see that $iZ_{L}$ belongs to the generated Lie algebra as well. Since $\{iX_{L},iY_{L},iZ_{L}\}$ is a basis of $\mathfrak{su}(2)$, it follows that
$\mathrm{Lie}\langle iX_{L},iY_{L}\rangle=\mathfrak{su}(2).$ (64)
Because the connected Lie group with Lie algebra $\mathfrak{su}(2)$ is $SU(2)$, the connected subgroup generated by the two tunable one–parameter families $\{P_{L}(\alpha)\}_{\alpha\in\mathbb{R}}$ and $\{B_{L}(\beta)\}_{\beta\in\mathbb{R}}$ is all of $SU(2)$ acting on $\mathcal{C}$. This proves that the SPE code supports arbitrary logical single–qubit unitaries using only the physically available phase shifter and two–mode coupler, once their action is represented in the $U(\Lambda_{\infty})$ eigenbasis of the code.
5 Entangling Two Single Photon Entanglement Qubits
In this section we construct a logical entangling gate using a coherent single–qubit $SU(2)$ control on the code space $\mathcal{C}$ (see section  4) and a two–photon polarisation parity measurement, which is a two–outcome CPTP instrument on the underlying optical degrees of freedom.
In order to establish the result we will proceed as follows:
  1. 1.
    We write the parity measurement as a quantum instrument defined on two SPE qubits state space in Kraus form [7] and explicit it in the logical basis.
  2. 2.
    We show that the the two qubits observables of the previous measurement performed on two input qubits $C$ and $D$:
    $S_{1}=X_{L}^{(C)}Z_{L}^{(D)},\ S_{2}=Z_{L}^{(C)}X_{L}^{(D)}$ (65)
    have a unique common +1+1 eigenstate and is maximally entangled.
  3. 3.
    We derive the Choi–Jamiołkowski [8], [9] resource state associated with the logical $CZ_{L}$ channel and show it is the unique common +1+1 eigenstate of the commuting stabilisers $\widetilde{S}_{i}$, $i=1,2,3,4$ representing parity polarization measurements on four SPE qubits.
  4. 4.
    We obtain a logical $CNOT_{L}$ gate up to Pauli products in two SPE qubits logical state space via a teleportation protocol built from the Choi representation defined at previous point.
5.1 Parity Polarisation Measurement as a CPTP Instrument
On two photons $A,B$ we consider the physical polarisation parity observable
$M_{ZZ}^{(\mathrm{pol})}:=I_{\rm dir}^{(A)}\otimes\sigma_{z}^{(\mathrm{pol},A)}\otimes I_{\rm dir}^{(B)}\otimes\sigma_{z}^{(\mathrm{pol},B)},\qquad\sigma_{z}|H\rangle=|H\rangle,\ \sigma_{z}|V\rangle=-|V\rangle.$ (66)
The two outcomes $s=\pm 1$ are described by projectors
$\Pi_{s}^{(\mathrm{pol})}=\frac{1}{2}\Big(I+s\,\sigma_{z}^{(\mathrm{pol},A)}\sigma_{z}^{(\mathrm{pol},B)}\Big),$ (67)
and ideal parity readout is the two–outcome instrument
$\mathcal{E}_{s}(\rho)=\Pi_{s}^{(\mathrm{pol})}\,\rho\,\Pi_{s}^{(\mathrm{pol})},\qquad s\in\{+1,-1\}.$ (68)
Thus, conditioned on $s$, the map has a single Kraus operator $K_{s}=\Pi_{s}^{(\mathrm{pol})}$; and the unconditional map $\mathcal{E}=\mathcal{E}_{+}+\mathcal{E}_{-}$ is CPTP.
Recall the single–photon two–mode basis $|\Phi_{+H}\rangle:=|+\rangle|H\rangle,\ |\Phi_{-V}\rangle:=|-\rangle|V\rangle$ and the logical SPE code basis
$|0_{L}\rangle=\frac{|\Phi_{+H}\rangle+|\Phi_{-V}\rangle}{\sqrt{2}},\qquad|1_{L}\rangle=\frac{|\Phi_{+H}\rangle-|\Phi_{-V}\rangle}{\sqrt{2}}.$ (69)
Since $\sigma_{z}^{(\mathrm{pol})}|\Phi_{+H}\rangle=+|\Phi_{+H}\rangle$ and $\sigma_{z}^{(\mathrm{pol})}|\Phi_{-V}\rangle=-|\Phi_{-V}\rangle,$ one finds
$\sigma_{z}^{(\mathrm{pol})}|0_{L}\rangle=|1_{L}\rangle,\qquad\sigma_{z}^{(\mathrm{pol})}|1_{L}\rangle=|0_{L}\rangle,\qquad\Rightarrow\qquad\sigma_{z}^{(\mathrm{pol})}\big|_{\mathcal{C}}=X_{L}.$ (70)
Therefore, on $\mathcal{C}_{A}\otimes\mathcal{C}_{B}$,
$\Big(\sigma_{z}^{(\mathrm{pol},A)}\sigma_{z}^{(\mathrm{pol},B)}\Big)\Big|_{\mathcal{C}^{(2)}}=X_{L}^{(A)}X_{L}^{(B)},\qquad\Pi_{s}^{(\mathrm{pol})}\Big|_{\mathcal{C}^{(2)}}=\frac{1}{2}\Big(I+s\,X_{L}^{(A)}X_{L}^{(B)}\Big).$ (71)
Hence, in the ideal representation–preserving limit, the physical parity device realises a logical Pauli–product measurement of $X_{L}\otimes X_{L}$. Using available single–qubit $SU(2)$ rotations, the same primitive realises any Pauli–product measurement needed below (e.g. $Z_{L}Z_{L}$).
5.2 Existence of Entangled States From Parity Polarization Measurements
Consider two logical SPE qubits labelled $(C,D)$. Define the commuting Pauli products
$S_{1}:=X_{L}^{(C)}Z_{L}^{(D)},\qquad S_{2}:=Z_{L}^{(C)}X_{L}^{(D)}.$ (72)
Commutation can be verified by the reader.
We now determine the unique common +1+1 eigenstate explicitly. Write a general logical two–qubit state as
$|\psi\rangle=a|0_{L}0_{L}\rangle+b|0_{L}1_{L}\rangle+c|1_{L}0_{L}\rangle+d|1_{L}1_{L}\rangle.$
Imposing $S_{1}|\psi\rangle=|\psi\rangle$ and $S_{2}|\psi\rangle=|\psi\rangle$ gives
$c=a,\qquad d=-b,\qquad b=a,\qquad d=-c,$
hence $b=c=a$ and d=ad=-a. Up to normalisation, there is a unique solution:
$|C_{2}\rangle_{CD}=\frac{1}{2}\Big(|0_{L}0_{L}\rangle+|0_{L}1_{L}\rangle+|1_{L}0_{L}\rangle-|1_{L}1_{L}\rangle\Big)_{CD}.$ (73)
This state is maximally entangled. Indeed, the coefficient matrix in the logical basis is $M=\frac{1}{2}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$, so $MM^{\dagger}=\mathbf{1}/2$ and the reduced state is maximally mixed.
Operationally, $|C_{2}\rangle$ is prepared inside $\mathcal{C}^{(2)}$ by measuring $S_{1}$ and $S_{2}$. If one obtains the (1)(-1) outcome for either stabiliser, the resulting state differs from (73) by a known local Pauli byproduct (tracked classically or corrected).
5.3 Controlled-$NOT$ Gate Construction
In this subsection we consider the above defined parity polarization measurements on four SPE qubits, represent it in logical basis and show it has a unique eigenstate with eigenvalue +1. We call such state $|R_{CZ}\rangle$. We then show that $|R_{CZ}\rangle$ can be used to construct a $CNOT_{L}$ gate defined on two logical qubits up to Pauli products exploiting a suitable teleportation protocol.
Introduce four logical qubits $C,C^{\prime},D,D^{\prime}$ and consider the following set of measurements:
$\widetilde{S}_{1}$ $=Z_{L}^{(C)}Z_{L}^{(C^{\prime})},$ (74)
$\widetilde{S}_{2}$ $=Z_{L}^{(D)}Z_{L}^{(D^{\prime})},$
$\widetilde{S}_{3}$ $=X_{L}^{(C)}X_{L}^{(C^{\prime})}Z_{L}^{(D)},$
$\widetilde{S}_{4}$ $=X_{L}^{(D)}X_{L}^{(D^{\prime})}Z_{L}^{(C)}.$
In the next subsection we are going to construct explicitly these three qubits measurements via ancilla assisted reduction of three qubits Pauli tensor products measurements to two qubits polarization parity measurements plus ancillas.
5.3.1 From Three Qubits Measurements to Polarization Parity Measurements
The stabilisers $\widetilde{S}_{3},\widetilde{S}_{4}$ in (74) have involve measurements on three qubits. Here we show explicitly how to implement a generic projective measurement on three qubits using: (i) the two-qubit parity measurement primitive (which realises $X_{L}\otimes X_{L}$ in the logical basis, cf. (71)), and (ii) single-qubit $SU(2)$ control on $\mathcal{C}$ (Sec. 4.2).
For any pair of logical qubits $(q,r)$, measuring $Z_{L}^{(q)}Z_{L}^{(r)}$ is equivalent to measuring $X_{L}^{(q)}X_{L}^{(r)}$ after a basis change:
$\Pi^{(ZZ)}_{\pm}(q,r):=\frac{1}{2}\Big(I\pm Z_{L}^{(q)}Z_{L}^{(r)}\Big)=(H_{L}^{(q)}\otimes H_{L}^{(r)})\;\frac{1}{2}\Big(I\pm X_{L}^{(q)}X_{L}^{(r)}\Big)\;(H_{L}^{(q)}\otimes H_{L}^{(r)}),$ (75)
since $H_{L}X_{L}H_{L}=Z_{L}$.
To measure $\widetilde{S}_{3}=X_{L}^{(C)}X_{L}^{(C^{\prime})}Z_{L}^{(D)}$, we equivalently measure
$S^{(3)}_{ZZZ}:=Z_{L}^{(C)}Z_{L}^{(C^{\prime})}Z_{L}^{(D)}$ (76)
after applying $H_{L}$ on the two qubits $(C,C^{\prime})$ (or, equivalently, by interpreting the measurement in a rotated basis). Indeed
$(H_{L}^{(C)}\otimes H_{L}^{(C^{\prime})}\otimes I^{(D)})\;S^{(3)}_{ZZZ}\;(H_{L}^{(C)}\otimes H_{L}^{(C^{\prime})}\otimes I^{(D)})=X_{L}^{(C)}X_{L}^{(C^{\prime})}Z_{L}^{(D)}=\widetilde{S}_{3}.$ (77)
Thus it suffices to give a constructive measurement of the three qubits product measurement $ZZZ$ (76) using only two-qubit parity measurements.
Introduce a helper logical ancilla $g$ prepared in the computational basis state $|0_{L}\rangle_{g}$ satisfying $Z_{L}^{(g)}|0_{L}\rangle_{g}=|0_{L}\rangle_{g}$, hence $Z_{L}^{(g)}=+1$ on the prepared state. Define the even-weight operator
$S^{(4)}_{ZZZZ}:=Z_{L}^{(C)}Z_{L}^{(C^{\prime})}Z_{L}^{(D)}Z_{L}^{(g)}.$ (78)
Because $Z_{L}^{(g)}=+1$ is fixed by preparation, measuring $S^{(4)}_{ZZZZ}$ is equivalent to measuring $S^{(3)}_{ZZZ}$.
Introduce a probe logical ancilla $p$ prepared in the $X_{L}$-eigenstate
$|+\rangle_{L}^{(p)}=\frac{|0_{L}\rangle_{p}+|1_{L}\rangle_{p}}{\sqrt{2}},\qquad X_{L}^{(p)}|+\rangle_{L}^{(p)}=|+\rangle_{L}^{(p)}.$ (79)
Now perform the four commuting two-qubit parity measurements
$Z_{L}^{(C)}Z_{L}^{(p)},\qquad Z_{L}^{(C^{\prime})}Z_{L}^{(p)},\qquad Z_{L}^{(D)}Z_{L}^{(p)},\qquad Z_{L}^{(g)}Z_{L}^{(p)},$ (80)
implemented physically by the available $X_{L}X_{L}$ parity primitive using (75). Let the outcomes be $s_{C},s_{C^{\prime}},s_{D},s_{g}\in\{\pm 1\}$, so that the post-measurement state lies in the joint eigenspace
$Z_{L}^{(C)}Z_{L}^{(p)}=s_{C},\quad Z_{L}^{(C^{\prime})}Z_{L}^{(p)}=s_{C^{\prime}},\quad Z_{L}^{(D)}Z_{L}^{(p)}=s_{D},\quad Z_{L}^{(g)}Z_{L}^{(p)}=s_{g}.$ (81)
Multiplying the four relations in (81) gives
$\big(Z_{L}^{(C)}Z_{L}^{(C^{\prime})}Z_{L}^{(D)}Z_{L}^{(g)}\big)\;\big(Z_{L}^{(p)}\big)^{4}=s_{C}\,s_{C^{\prime}}\,s_{D}\,s_{g}.$ (82)
Since $(Z_{L}^{(p)})^{2}=I$, one has $(Z_{L}^{(p)})^{4}=I$, hence the product outcome
$s:=s_{C}\,s_{C^{\prime}}\,s_{D}\,s_{g}\in\{\pm 1\}$ (83)
is exactly the measured eigenvalue of the even-weight operator (78):
$S^{(4)}_{ZZZZ}=s.$ (84)
Because $Z_{L}^{(g)}=+1$ by preparation, (84) is equivalent to
$S^{(3)}_{ZZZ}=Z_{L}^{(C)}Z_{L}^{(C^{\prime})}Z_{L}^{(D)}=s.$ (85)
This implements a refinement of the $ZZZ$ projective measurement; in particular the post-measurement state is conditioned on the full tuple $(s_{C},s_{C}^{\prime},s_{D},s_{g})$, and therefore lies in the $ZZZ$ eigenspace with eigenvalue $s=s_{C}s_{C^{\prime}}s_{D}s_{g}$. Finally, undoing the basis change in (77) (or, equivalently, interpreting the whole procedure in the rotated basis) yields a projective measurement of $\widetilde{S}_{3}$ with outcome $s$:
$\widetilde{S}_{3}=X_{L}^{(C)}X_{L}^{(C^{\prime})}Z_{L}^{(D)}=s.$ (86)
The helper $g$ and probe $p$ may then be discarded (e.g. by measuring them in a fixed basis), since the logical information relevant to the stabiliser constraints is now stored on the data registers.
To measure $\widetilde{S}_{4}=X_{L}^{(D)}X_{L}^{(D^{\prime})}Z_{L}^{(C)}$, apply the same steps with the replacements $(C,C^{\prime},D)\mapsto(D,D^{\prime},C)$. In both cases, only the two-qubit parity primitive and single-qubit basis changes are used.
5.3.2 Parity Polarization Measurements and Controlled-$Z_{L}$ Gate
Each $\widetilde{S}_{i}$ is Hermitian with eigenvalues $\pm 1$, and one checks directly that $\widetilde{S}_{i},\widetilde{S}_{j}]=0$ for all $i,j$. Therefore the four observables admit a simultaneous eigenbasis. The eigenvector is unique because the four commuting stabilisers are independent and generate a stabiliser group of size $2^{4}$ acting on $4$ qubits, hence they fix a one–dimensional subspace in $\mathcal{C}^{\otimes 4}$. In order to find the form of $|R_{CZ}\rangle$ in logical basis note that:
$\widetilde{S}_{1}=+1\ \Rightarrow\ Z_{C}=Z_{C^{\prime}},\qquad\widetilde{S}_{2}=+1\ \Rightarrow\ Z_{D}=Z_{D^{\prime}}.$
Thus, in logical basis, the support of $|R_{CZ}\rangle$ lies in the span of $\{|00\rangle_{CC^{\prime}},|11\rangle_{CC^{\prime}}\}\otimes\{|00\rangle_{DD^{\prime}},|11\rangle_{DD^{\prime}}\}.$ Write
$|R_{CZ}\rangle=a\,|00\rangle_{CC^{\prime}}|00\rangle_{DD^{\prime}}+b\,|00\rangle_{CC^{\prime}}|11\rangle_{DD^{\prime}}+c\,|11\rangle_{CC^{\prime}}|00\rangle_{DD^{\prime}}+d\,|11\rangle_{CC^{\prime}}|11\rangle_{DD^{\prime}}.$ (87)
Now impose $\widetilde{S}_{3}=+1$ and $\widetilde{S}_{4}=+1$. Using that $X_{C}X_{C^{\prime}}$ swaps $|00\rangle_{CC^{\prime}}\leftrightarrow|11\rangle_{CC^{\prime}}$ and that $Z_{D}$ acts as +1+1 on $|00\rangle_{DD^{\prime}}$ and 1-1 on $|11\rangle_{DD^{\prime}}$, the constraint $\widetilde{S}_{3}|R_{CZ}\rangle=|R_{CZ}\rangle$ gives
$c=a,\qquad d=-b.$ (88)
Similarly, since $X_{D}X_{D^{\prime}}$ swaps $|00\rangle_{DD^{\prime}}\leftrightarrow|11\rangle_{DD^{\prime}}$ and $Z_{C}$ acts as +1+1 on $|00\rangle_{CC^{\prime}}$ and 1-1 on $|11\rangle_{CC^{\prime}}$, the constraint $\widetilde{S}_{4}|R_{CZ}\rangle=|R_{CZ}\rangle$ gives
$b=a,\qquad d=-c.$ (89)
Combining (88) and (89) yields $b=c=a$ and d=a.d=-a. After normalisation we obtain the explicit expression of $|R_{CZ}\rangle_{CC^{\prime}DD^{\prime}}$
$|R_{CZ}\rangle_{CC^{\prime}DD^{\prime}}=\frac{1}{2}\Big(|00\rangle_{CC^{\prime}}|00\rangle_{DD^{\prime}}+|00\rangle_{CC^{\prime}}|11\rangle_{DD^{\prime}}+|11\rangle_{CC^{\prime}}|00\rangle_{DD^{\prime}}-|11\rangle_{CC^{\prime}}|11\rangle_{DD^{\prime}}\Big).$ (90)
We now show that (90) coincides with the Choi–Jamiołkowski state of the unitary Controlled $Z_{L}$ quantum gate with input registers $(C^{\prime},D^{\prime})$ and output registers $(C,D)$.
The action of a unitary $CZ_{L}$ gate on the logical basis two qubits basis is:
$CZ_{L}|x_{L}\rangle|y_{L}\rangle=(-1)^{xy}|x_{L}\rangle|y_{L}\rangle$ (91)
Define the logical EPR pairs
$|\Phi^{+}\rangle_{CC^{\prime}}=\frac{|00\rangle_{CC^{\prime}}+|11\rangle_{CC^{\prime}}}{\sqrt{2}},\qquad|\Phi^{+}\rangle_{DD^{\prime}}=\frac{|00\rangle_{DD^{\prime}}+|11\rangle_{DD^{\prime}}}{\sqrt{2}}.$ (92)
Applying $CZ_{L}$ on the output qubits $(C,D)$ to the product of EPR pairs gives the Choi representation of $CZ_{L}$:
$(I_{C^{\prime}D^{\prime}}\otimes CZ_{L}^{(C,D)})\,\big(|\Phi^{+}\rangle_{CC^{\prime}}\otimes|\Phi^{+}\rangle_{DD^{\prime}}\big)=\frac{1}{2}\Big(|00\rangle_{CC^{\prime}}|00\rangle_{DD^{\prime}}+|00\rangle_{CC^{\prime}}|11\rangle_{DD^{\prime}}+|11\rangle_{CC^{\prime}}|00\rangle_{DD^{\prime}}-|11\rangle_{CC^{\prime}}|11\rangle_{DD^{\prime}}\Big),$ (93)
which coincides exactly with (90). Hence
$|R_{CZ}\rangle_{CC^{\prime}DD^{\prime}}=(I_{C^{\prime}D^{\prime}}\otimes CZ_{L}^{(C,D)})\,\big(|\Phi^{+}\rangle_{CC^{\prime}}\otimes|\Phi^{+}\rangle_{DD^{\prime}}\big).$ (94)
which shows equivalence between $|R_{CZ}\rangle_{CC^{\prime}DD^{\prime}}$ and the Choi representation of $CZ_{L}$ gate on two input qubits.
5.3.3 Controlled NOT Gate via Teleportation
Exploiting the above construction we are now going to define a teleportation protocol that implements a $CNOT_{L}$ on an arbitrary two qubits state $|\psi\rangle_{AB}$ teleporting it on two output qubits state $|\text{out}\rangle_{CD}$.
Let $(A,B)$ be the two input logical SPE qubits in an arbitrary state $|\psi\rangle_{AB}$. Assume we have prepared the four–qubit resource $|R_{CZ}\rangle_{CC^{\prime}DD^{\prime}}$.
We now perform logical Bell measurements on the pairs $(A,C^{\prime})$ and $(B,D^{\prime})$. A logical Bell measurement on a pair $(q,r)$ is equivalent to measuring the commuting observables
$Z_{L}^{(q)}Z_{L}^{(r)}\quad\text{and}\quad X_{L}^{(q)}X_{L}^{(r)}.$ (95)
In our architecture, $X_{L}X_{L}$ is directly measured by polarisation parity (see  (71)), while $Z_{L}Z_{L}$ is obtained by conjugation with $H_{L}\otimes H_{L}$ (available from single–qubit $SU(2)$ control).
Encode the Bell outcomes on $(A,C^{\prime})$ as bits $(x,y)\in\{0,1\}^{2}$:
$Z_{L}^{(A)}Z_{L}^{(C^{\prime})}=(-1)^{x},\qquad X_{L}^{(A)}X_{L}^{(C^{\prime})}=(-1)^{y},$ (96)
and similarly on $(B,D^{\prime})$ as $(l,m)\in\{0,1\}^{2}$:
$Z_{L}^{(B)}Z_{L}^{(D^{\prime})}=(-1)^{l},\qquad X_{L}^{(B)}X_{L}^{(D^{\prime})}=(-1)^{m}.$ (97)
After these Bell measurements, the measured qubits are $A,C^{\prime},B,D^{\prime}$, hence they are no longer part of the quantum output. The only remaining quantum system is precisely the pair $(C,D)$.
The teleportation identity implies that the postmeasurement state on $(C,D)$ is
$|\mathrm{out}\rangle_{CD}=\Big(X_{L}^{y}Z_{L}^{x}\Big)^{(C)}\Big(X_{L}^{m}Z_{L}^{l}\Big)^{(D)}\;CZ_{L}\;|\psi\rangle_{AB}.$ (98)
Here $CZ_{L}|\psi\rangle_{AB}$ is understood as the same logical vector, with the information teleported from registers $(A,B)$ onto $(C,D)$, and acted on by $CZ_{L}$. This equation implements $CZ_{L}$ up to Pauli byproducts: the byproducts are known local Pauli operators determined by the classical outcome bits $(x,y,l,m)$, and are therefore exactly correctable by a final single–qubit correction or by tracking a classical Pauli frame.
Finally,
$CNOT_{L}=\big(\mathbf{1}\otimes H_{L}\big)\;CZ_{L}\;\big(\mathbf{1}\otimes H_{L}\big),$ (99)
so the parity–based construction yields a logical $CNOT_{L}$ gate on the SPE code.
5.4 Circuit-model universality
Since we can implement arbitrary single–qubit unitaries $SU(2)$ on $\mathcal{C}$ (section  4.2) and we can implement an entangling two–qubit gate $CNOT_{L}$ (section  5.3), the SPE model is universal for quantum computation in the standard circuit sense.
6 Conclusion
We have studied the massless sector of the extended Poincaré group $\mathcal{P}_{\mathrm{ext}}$ introduced in [1]. Because the lightlike little group is enlarged to $ISO(2)\rtimes\mathbb{Z}_{2}$, the corresponding massless UIRs necessarily carry a binary internal structure: the representation space decomposes as a direct sum of forward and backward Wigner sectors, $\mathcal{H}_{\oplus}=\mathcal{H}_{\mathrm{fwd}}\oplus\mathcal{H}_{\mathrm{bwd}}$, and the discrete involutive element acts as a sector–exchange symmetry. We showed that this binary structure admits an operationally faithful embedding into a bipartite system $\mathcal{H}\otimes\mathbb{C}^{2}$: there exist a sector isometry $V$ and a natural *–homomorphism $\iota$ of observable algebras such that expectation values of all local observables $A\otimes B$ coincide with those of the corresponding operators $\iota(A\otimes B)$ on $\mathcal{H}_{\oplus}$. In a two–mode single–photon restriction, this reduces to ordinary two–qubit Bell-like states and Pauli correlations.
We then connected the abstract label $\varepsilon=\pm 1$ to an experimentally accessible signature. For a fixed propagation axis $\hat{n}$ the two sectors can be realised as counterpropagating optical modes, while a two-level internal degree of freedom (e.g. polarisation) encodes the binary label. In this setting the correlation observable $\sigma_{x}^{(\mathrm{dir})}\otimes\sigma_{x}^{(\mathrm{pol})}$ has expectation value $\varepsilon$ on the $U(\Lambda_{\infty})$–eigenstates, providing a direct tomographic handle on the sector degree of freedom and a concrete route to empirical constraints on the extended massless representation.
Finally, we used the same single–photon two–mode restriction as a computational code. The two involution eigenstates define a logical qubit (the single photon entanglement qubit) whose pure states form a Bloch sphere and whose logical generators close $\mathfrak{su}(2)$; physically, a relative phase shifter and a two–mode coupler generate all logical single–qubit unitaries. On two such qubits, a polarisation parity measurement restricts to a logical Pauli–product measurement on the code space. From this primitive we constructed a Choi resource state for $CZ_{L}$ as a stabiliser state and implemented $CZ_{L}$ (hence $CNOT_{L}$) on arbitrary inputs via a gate–teleportation protocol, up to Pauli byproducts. Together, arbitrary single–qubit $SU(2)$ control and an entangling two–qubit gate establish circuit–model universality of the SPE scheme. In the computational model presented, the logical representation space coincides with the physical Hilbert space of the Poincarè group, and logical operations are equivalent to idealized physical transformations. This identification has the consequence that quantum entanglement appears as a structural feature of representation theory rather than an externally engineered resource and the leakage of resources in a computation from computational space to physical degrees of freedom not relevant to the computation is not part of the definition of the computation itself but is in principle restricted to noise and imperfections of optical gates implementing logical operations.
This could be a conceptual breakthrough in the developments of architectures for scalable quantum computers.